### v1

# sharp bottom spectrum and scalar curvature rigidity

Abstract.
We prove a sharp upper bound for the bottom spectrum of Laplacian on geometrically contractible manifolds with scalar curvature lower bound, and characterize the distribution of scalar curvature when equality holds. Moreover, we prove a scalar curvature rigidity theorem if the manifold is the universal cover of a closed hyperbolic manifold.
The first author is partially supported by NSF 1952693.
1. Introduction
Let $(X^{n},g)$ be a connected, complete, noncompact Riemannian manifold, and $\Delta$ the corresponding Beltrami Laplacian on $(X^{n},g)$ defined by
$\Delta f=\sum_{i=1}^{n}\left(\nabla_{e_{i}}\nabla_{e_{j}}-\nabla_{\nabla_{e_{i}}e_{j}}\right)f=\frac{1}{\sqrt{\det(g)}}\frac{\partial}{\partial x^{i}}\left(\sqrt{\det(g)}g^{ij}\frac{\partial f}{\partial x^{j}}\right).$
The $L^{2}$-bottom spectrum of $\Delta$ is defined by
$\lambda_{1}(X,g)=\inf\left\{\frac{\int_{X}|\nabla f|^{2}}{\int_{X}f^{2}}:\ f\in C_{c}^{\infty}(M),\ f\neq 0\right\}$ (1.1)
Recall that the classical comparison theorem in [Cheng_eigenvalue] implies that, if $(X^{n},g)$ satisfies $\mathrm{Ric}_{g}\geq-(n-1)$, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}.$
A natural question is to generalize this question to scalar curvature lower bound. Recently, Munteanu–Wang in [MunteanuWang, Theorem 1.1] generalize the sharp bottom spectrum estimate on three-dimensional manifolds with scalar curvature lower bound. If $(X^{3},g)$ is a complete Riemannian manifold with scalar curvature $\mathrm{Sc}_{g}\geq-6$ and satisfies either one of the following properties:
  • the second homology group $H_{2}(X,\mathbb{Z})$ contains no spherical class, or
  • $X$ has finitely many ends and finite first Betti number $b_{1}(X)<\infty$,
then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}.$
In this paper, we further study the bottom spectrum under scalar curvature lower bound using the Dirac operator method and the higher index theory. Higher index theory [BaumConnesHigson, MR0996437, Roe] is a far-reaching generalization of the classical Fredholm index, particularly for non-compact manifolds, and is a more refined index theory than the theorems of Atiyah–Singer. We will later give a brief introduction for the basic setups for higher indices in Section 2.
Let us recall that a Riemannian manifold is said to be geometrically contractible if there exists a function $R(r)\geq r$ for any $r\geq 0$ such that $B(p,r)$ is contractible in $B(p,R(r))$ for any $p\in M$. Note that the universal cover of any aspherical manifold is geometrically contractible (see [wang2024fillingradiusquantitativektheory, Example 2.6]). A complete Riemannian manifold is said to have bounded geometry if the sectional curvature and its derivatives are uniformly bounded and injective radius has uniformly lower bound, i.e., $|\nabla^{\alpha}\sec_{g}|\leq K_{\alpha},\ \operatorname{Inj}(M)\geq i>0$ for any multi-index $\alpha$ and constants $K_{\alpha},i>0$. The first main theorem of this paper is as follows.
Theorem 1.1.
Suppose that $(X^{n},g)$ is complete, geometrically contractible Riemannian manifold with bounded geometry and scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)\kappa$ for some constant $\kappa\geq 0$. If $(X^{n},g)$ satisfies the coarse Novikov conjecture, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}\kappa.$ (1.2)
Moreover, if $\lambda_{1}(X,g)=\frac{n-1}{4n}\kappa$, then for any $\delta>0$, the set
$\{p\in X:\mathrm{Sc}_{g}(p)\geq-\kappa+\delta\}$
is not a net of $({X},g)$.
The precise statement of the coarse Novikov conjecture is in Conjecture 2.6. Recall that a subset $S$ in $X$ is is said to be a net of $X$ if there exists $r>0$ such that $N_{r}(S)=X$, where $N_{r}(S)=\{x\in X:\mathrm{dist}(x,S)<r\}$. Theorem 1.1 is a geometric version of a more general theorem presented in Section 4. We emphasize that the upper bound $\frac{(n-1)^{2}}{4}\kappa$ is sharp, as the standard hyperbolic space $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$ has scalar curvature $-n(n-1)$ and $\lambda_{1}(\mathbb{H}^{n},g_{\mathbb{H}^{n}})=\frac{(n-1)^{2}}{4}$.
Classically, Li-Wang make use of the harmonic function theory to study the sharpness of the upper bound on the bottom spectrum and the splitting rigidity in [Li_Wang_positive_spectrum_1, Li_Wang_positive_spectrum_2] if the Ricci curvature $\mathrm{Ric}_{g}\geq-(n-1)$. As pointed out in [MunteanuWang], it is impossible to generalize the classical sharp bottom spectrum result to any complete Riemannian manifold with only scalar curvature assumption. Essentially, our philosophy for proving Theorem 1.1 is based on the Novikov conjecture. Moreover, the method, using higher index theory as a starting point, has a broader application since it is easier to achieve non-vanishing higher index than non-vanishing Fredholm index or $L^{2}$-index used in [Davaux] and has no restriction on the parity of the dimension. The rigidity result [Davaux, Theorem C] is modeled on the flat space due to the assumption of nonzero $\hat{A}$-genus. See the survey related to Novikov conjecture in [Xie_Yu_higher_invariants] or a list at the end of the Section 2 in this paper.
Furthermore, the net characterization on general complete noncompact manifold, as equality holds, is possibly the best expectation and cannot be further improved in general, which is supported by the following example inspired by [MunteanuWang, Theorem 1.4 & Example 1.5].
Example 1.2.
Let $X=\mathbb{R}^{n-1}\times\mathbb{R}$ be the complete manifold equipped with the metric
$g=dt^{2}+\cosh^{\frac{2}{a}}(at)g_{\mathbb{R}^{n-1}},$
where $(n-1)/2\leq a<n/2$. Note that $(X,g)$ is a geometrically contractible manifold with bounded geometry and satisfies the coarse Novikov conjecture [willett2020higher, Chapter 7]. A direct calculation shows that
  1. (1)
    $\mathrm{Sc}_{g}=-n(n-1)+(n-1)\frac{n-2a}{\cosh^{2}(t)}>-n(n-1)$;
  2. (2)
    $\lambda_{1}(X,g)=\frac{(n-1)^{2}}{4}$ since
    $\Delta(\cosh^{-\frac{n-1}{2a}}(at))$
    $=$ $\cosh^{-\frac{n-1}{a}}(at)\partial_{t}(\cosh^{\frac{n-1}{a}}(at)\partial_{t}(\cosh^{-\frac{n-1}{2a}}(at)))$
    $=$ $\cosh^{-\frac{n-1}{a}}(at)\partial_{t}(-\frac{n-1}{2}\cosh^{\frac{n-1}{2a}-1}(at)\sinh(at))$
    $=$ $-\frac{n-1}{2}\big{(}\frac{n-1}{2}-a\big{)}\cosh^{-\frac{n-1}{2a}-2}(at)\sinh^{2}(at)-a\frac{n-1}{2}\cosh^{-\frac{n-1}{2a}}(at)$
    $\leq$ $-\frac{(n-1)^{2}}{4}\cosh^{-\frac{n-1}{2a}}(at).$
Moreover, if the Riemannian manifold is the universal cover of a closed manifold, we further obtain a scalar curvature rigidity. Here, we only state the theorem in case of closed hyperbolic manifold.
Theorem 1.3.
Suppose that $M^{n}$ is a closed hyperbolic manifold. If $(M^{n},g)$ has scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)\kappa$ for some constant $\kappa\geq 0$, then
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{(n-1)^{2}}{4}\kappa,$ (1.3)
where $(\widetilde{M},\widetilde{g})$ is the Riemannian universal cover of $(M,g)$. Moreover, if the equality holds, then $Sc_{g}=-n(n-1)\kappa$.
A more general result is given Section 4, showing that the conclusion also holds for closed essential manifolds whose fundamental groups satisfy the strong Novikov conjecture 2.7.
It is known that closed hyperbolic manifold admits no complete Riemannian metric with non-negative scalar curvature. Note that Theorem 1.3 can be rewritten as
$\inf_{p\in M}(-\mathrm{Sc}_{g}(p))\leq\frac{4n}{n-1}\lambda_{1}(\widetilde{M},\widetilde{g}).$
Hence, it can be viewed as a quantitative version of the nonexistence of complete Riemannian metric with non-negative scalar curvature on hyperbolic manifold. A further question in this direction is to find an optimal upper bound on $\inf_{p\in M}(-\mathrm{Sc}_{g}(p))$ in terms of the group complexity invariant of the fundamental group for hyperbolic manifold or aspherical manifold, and the $\operatorname{\mathrm{vol}}(M)$. Moreover, it is interesting to study the relations between the bottom spectrum $\lambda_{1}$ and the macroscopic scalar curvature, the seminal works [Guth_volumes_of_balls, Roman_volume, B_K_macroscopic, Sabine_Roman_volume, Hannah_Kei_Area, alpert2024growthuniversalcoverlarge] provide series of relations between the macroscopic curvature and other geometric and topological invariants. Finally, we note that Wang [Wang_eigenvalue] prove that, if $\mathrm{Ric}_{g}\geq-(n-1)$, the $\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{(n-1)^{2}}{4}$. In particular, the equality holds if and only if $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$. Theorem 1.3 directly leads to a further rigidity conjecture as follows.
Conjecture 1.4.
Suppose that $M^{n}$ is a closed hyperbolic manifold. If $(M^{n},g)$ has scalar curvature $\mathrm{Sc}_{g}=-n(n-1)$ and
$\lambda_{1}(\widetilde{M},\widetilde{g})=\frac{(n-1)^{2}}{4},$
then $(\widetilde{M},\widetilde{g})$ isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$.
Note that Conjecture 1.4 has been confirmed in [MunteanuWang] for $n=3$. Finally, Theorem 1.3 is connected with other problems as follows. It is known that for closed Riemannian manifold,
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{h^{2}_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})}{4}$
with $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})$ the volume entropy. Ledrappier–Wang and Liu in [Fran_Wang_entropy_rigidity, Liu_volume_entropy] prove that, if $(M^{n},g)$ is a closed Riemannian manifold with $\mathrm{Ric}_{g}\geq-(n-1)$, then $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\leq n-1$, and the equality holds if and only if $(\widetilde{M},\widetilde{g})$ is isometric to the standard hyperbolic manifold $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$. However, Kazaras–Song–Xu in [kazaras2024scalarcurvaturevolumeentropy] prove that the volume entropy $h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\leq n-1$ does not hold for a closed hyperbolic manifold $(M^{n},g)$ with scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)$. Additionally, it’s known that there exists a constant $c_{n}$ such that
$\big{(}h_{\operatorname{\mathrm{vol}}}(\widetilde{M},\widetilde{g})\big{)}^{n}\operatorname{\mathrm{vol}}(M)\geq c_{n}\|M\|$
for any closed Riemannian manifold by Gromov in [Gromov_volume], and also a sharp constant $c_{n}$ has been obtained for the locally symmetric spaces due to Besson–Courtois–Gallot in [BCG_entropy, BCG_minimal_entropy]. Motivated by this, we propose the following question.
Problem 1.5.
Suppose that $(M^{n},g)$ is a closed (hyperbolic) manifold.
  • Study the invariant
    $\big{(}\lambda_{1}(\widetilde{M},\widetilde{g})\big{)}^{\frac{n}{2}}vol(M,g).$ (1.4)
  • We expect that there exists a constant $c_{n}$ such that
    $\|M\|\leq c_{n}\big{(}\lambda_{1}(\widetilde{M},\widetilde{g})\big{)}^{\frac{n}{2}}vol(M,g).$ (1.5)
    Here, $\|M\|$ is the simplicial volume of $M$(see [Gromov_volume] for the definition of simplicial volume).
Problem 1.5 holds trivially when $\pi_{1}(M)$ is amenable, where $\|M\|=\lambda_{1}(\widetilde{M},\widetilde{g})=0$. Moreover, Mohsen studies the first question in [Olivier_bottom_spectrum_hyperbolic, Theorem 1] and proves that the hyperbolic metric on $M$ is a saddle point of $\lambda_{1}(\widetilde{M})$. More precisely, the bottom spectrum is maximal among the conformal metrics of same volume, and minimal in its Ebin class. Problem 1.5 is connected with Gromov simplicial volume conjecture for general closed Riemannian manifold (see [Gromov_101, Section 26]) and Schoen conjecture on closed hyperbolic manifold (see [Schoen_variational_theory]), and macroscopic scalar curvature is a stronger concept than scalar curvature and many progresses has been made(see [Guth_volumes_of_balls, alpert2024growthuniversalcoverlarge, Sabine_Roman_volume] and the literature therein) in this direction. Hence, it is worthwhile of studying Problem 1.5 for its own interest and applications.
Outline
This paper is structured as follows: In Section 2, we provide the necessary background on higher index theory. In Section 3, we establish the Kato inequality for harmonic spinors, which is a standard result. We give a elementary proof and a slightly more general version for non-compact manifolds. In Section 4, we prove the main theorems presented in this paper and derive several related corollaries. Finally, in Section 5, to establish a net characterization for complete noncompact manifolds and demonstrate scalar curvature rigidity for cocompact manifolds, we prove a unique continuation theorem on complete Riemannian manifolds.
Acknowledgement
We would like to thank Shiqi Liu, Yuguang Shi, Jiaping Wang, Xiaodong Wang, Zhizhang Xie and Guoliang Yu for their interest and discussion on this topic.
2. Preliminaries on higher index theory
In this section, we review the construction of the geometric $C^{*}$-algebras and the higher index. The readers may refer to [willett2020higher] for more details.
2.1. Roe algebras and localization algebras
We will first review the definitions of some geometric $C^{*}$-algebras.
Suppose that $X$ is a proper metric space, i.e., every closed ball is compact. Let $\Gamma$ be a discrete group acting on $X$ by isometry. In the following, we only consider the cases where either $\Gamma$ is trivial, or $\Gamma$ acts properly and cocompactly. Let $C_{0}(X)$ be the $C^{*}$-algebra consisting of all complex-valued continuous functions on $X$ that vanish at infinity. A $\Gamma$-$X$-module is a separable Hilbert space $H_{X}$ equipped with a *-representation $\varphi$ of $C_{0}(X)$ and an action $\pi$ of $\Gamma$, which are compatible in the sense that
$\pi(\gamma)\big{(}\varphi(f)\xi\big{)}=\varphi(f^{\gamma})\big{(}\pi(\gamma)\xi\big{)},~{}\forall f\in C_{0}(X),~{}\gamma\in\Gamma,~{}\xi\in H_{X},$
where $f^{\gamma}(x)\coloneqq f(\gamma^{-1}x)$.
A $\Gamma$-$X$-module $H_{X}$ is called admissible if
  1. (1)
    $H_{X}$ is nondegenerate, namely the representation $\varphi$ is nondegenerate,
  2. (2)
    $H_{X}$ is standard, namely no nonzero function in $C_{0}(X)$ acts as a compact operator, and
  3. (3)
    for any $x\in X$, the stabilizer group $\Gamma_{x}$ acts on $H_{X}$ regularly, in the sense that the action is isomorphic to the action of $\Gamma_{x}$ on $l^{2}(\Gamma_{x})\otimes H$ for some infinite dimensional Hilbert space $H$.
For example, if $X$ is a $\Gamma$-cover of a closed manifold, then $L^{2}(X)$ is naturally a $\Gamma$-$X$-module.
Definition 2.1.
Let $H_{X}$ be an admissible $\Gamma$-$X$-module and $T$ is a bounded linear operator acting on $H_{X}$.
  1. (1)
    The propagation of $T$ is defined by
    $\mathrm{prop}(T)=\sup\{d(x,y)~{}|~{}(x,y)\in\text{supp}(T)\},$
    where $\text{supp}(T)$ is the complement (in $X\times X$) of the set of points $(x,y)\in X\times X$ such that there exists $f_{1},f_{2}\in C_{0}(X)$ such that $f_{1}Tf_{2}=0$ and $f_{1}(x)f_{2}(y)\neq 0$;
  2. (2)
    $T$ is said to be locally compact if both $fT$ and $Tf$ are compact for all $f\in C_{0}(X)$.
  3. (3)
    $T$ is said to be $\Gamma$-equivariant if $\gamma T=T\gamma$ for any $\gamma\in\Gamma$.
Definition 2.2.
Let $H_{X}$ be a standard nondegenerate $\Gamma$-$X$-module and $B(H_{X})$ the set of all bounded linear operators on $H_{X}$.
  1. (1)
    The equivariant $Roe$ algebra of $X$, denoted by $C^{*}(X)^{\Gamma}$, is the $C^{*}$-algebra generated by all locally compact, equivariant operators with finite propagation in $B(H_{X})$.
  2. (2)
    The equivariant localization algebra $C_{L}^{*}(X)^{\Gamma}$ is the $C^{*}$-algebra generated by all bounded and uniformly norm-continuous functions $f\colon[1,\infty)\to C^{*}(X)^{\Gamma}$ such that
    $\mathrm{prop}(f(t))<\infty\text{ and }\mathrm{prop}(f(t))\to 0\ \text{~{}as~{}}t\to\infty.$
The Roe algebras and localization algebras of $X$ are independent (up to isomorphisms) of the choice of nondegenerate standard $\Gamma$-$X$-modules $H_{X}$ [Yulocalization].
There is a natural evaluation map
$ev\colon C^{*}_{L}(X)^{\Gamma}\to C^{*}(X)$
induced by evaluating a path at $t=1$. The induced map $ev_{*}$ at the level of $K$-theory is also usually referred to as the index map or the assembly map.
We will omit $\Gamma$ if $\Gamma$ is trivial. In the case where $\Gamma$ acts on $X$ properly and cocompactly, we have that $C^{*}(X)^{\Gamma}\cong C^{*}_{r}(\Gamma)\otimes\mathcal{K}$, where $C^{*}_{r}(\Gamma)$ is the reduced group $C^{*}$-algebra of $\Gamma$ and $\mathcal{K}$ is the algebra of compact operators. In particular, we have $K_{*}(C^{*}(X)^{\Gamma})\cong K_{*}(C^{*}_{r}(\Gamma))$.
2.2. Higher index and local higher index
In this subsection, we will recall the definition of the higher index and local higher index for Dirac operators.
Let $\chi$ be a continuous function on $\mathbb{R}$. $\chi$ is said to be a normalizing function if it is non-decreasing, odd (i.e. $\chi(-x)=-\chi(x)$) and
$\lim_{x\to\pm\infty}\chi(x)=\pm 1.$
Suppose that $X$ is a complete spin manifold. Let $D$ be the associated Dirac operator on $X$ acting on the spinor bundle of $X$ and $\Gamma$ is a discrete group acting on $X$ isometrically. Moreover, let $H$ be the Hilbert space of the $L^{2}$-sections of the spinor bundle, which is an admissible $\Gamma$-$X$-module in the sense of Section 2.1. Let us first assume that $\dim X$ is even. In this case, the spinor bundle is naturally $\mathbb{Z}_{2}$-graded and the Dirac operator $D$ is an odd operator given by
$D=\begin{pmatrix}0&D_{+}\\ D_{-}&0\end{pmatrix}.$
Let $\chi$ be a normalizing function. Since $\chi$ is an odd function, we see that $\chi(t^{-1}D)$ is also a self-adjoint odd operator for any $t>0$ given by
$\chi(t^{-1}D)=\begin{pmatrix}0&U_{t,D}\\ V_{t,D}&0\end{pmatrix}.$ (2.1)
Now, we set
$W_{t,D}=\begin{pmatrix}1&U_{t,D}\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ -V_{t,D}&1\end{pmatrix}\begin{pmatrix}1&U_{t,D}\\ 0&1\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\ e_{1,1}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$
and
$\begin{split}P_{t,D}=&W_{t,D}e_{1,1}W_{t,D}^{-1}\\ =&\begin{pmatrix}1-(1-U_{t,D}V_{t,D})^{2}&(2-U_{t,D}V_{t,D})U_{t,D}(1-V_{t,D}U_{t,D})\\ V_{t,D}(1-U_{t,D}V_{t,D})&(1-V_{t,D}U_{t,D})^{2}\end{pmatrix}.\end{split}$ (2.2)
The path $(P_{t,D})_{t\in[1,+\infty)}$ defines an element in $M_{2}((C^{*}_{L}(X)^{\Gamma})^{+})$, and the difference $P_{t,D}-e_{1,1}$ lies in $M_{2}(C^{*}_{L}(X)^{\Gamma})$.
Definition 2.3.
If $X$ is a spin manifold of even dimension, then
  • the local higher index $\textup{Ind}_{L}(D)$ of $D$ is defined to be
    $\textup{Ind}_{L}(D)\coloneqq[P_{t,D}]-[e_{1,1}]\in K_{0}(C^{*}_{L}(X)^{\Gamma});$
  • the higher index $\textup{Ind}(D)$ of $D$ is defined to be
    $\textup{Ind}(D)\coloneqq[P_{1,D}]-[e_{1,1}]\in K_{0}(C^{*}(X)^{\Gamma}).$
The constructions of the (local) higher index for the odd dimensional spin manifold are as follows.
Definition 2.4.
If $X$ is a spin manifold of odd dimension, then
  • the local higher index $\textup{Ind}_{L}(D)$ of $D$ is defined to be
    $[e^{2\pi i\frac{\chi(t^{-1}D)+1}{2}}]\in K_{1}(C^{*}_{L}(X)^{\Gamma});$
  • the higher index $\textup{Ind}(D)$ of $D$ is defined to be
    $[e^{2\pi i\frac{\chi(D)+1}{2}}]\in K_{1}(C^{*}(X)^{\Gamma}).$
Note that the higher index and the local higher index are independent of the choices of normalizing functions. The $K$-theory $K_{\ast}(C^{*}_{L}(X)^{\Gamma})$ of the localization algebra $C^{*}_{L}(X)^{\Gamma}$ is naturally isomorphic to the $\Gamma$-equivariant $K$-homology of $X$. Under this isomorphism, the local higher index of $D$ coincides with the $K$-homology class of $D$. See [Yulocalization, QiaoRoe].
If $D$ is assumed to be an invertible operator on a spin manifold $X$, namely 00 is not in the spectrum of $D$, then the normalizing function $\chi$ can be the following function
$\chi(x)=\begin{cases}1&x\geq 0,\\ -1&x<0,\end{cases}$
which is continuous on the spectrum of $D$, and satisfies $\chi(D)^{2}=1$. Consquently, we reach that
  • if $X$ has even dimension, then $P_{1,D}=e_{1,1}$;
  • if $X$ has odd dimension, then
    $e^{2\pi i\frac{\chi(D)+1}{2}}=1.$
It follows that $\textup{Ind}(D)=0$. Hence, we conclude this section with the following proposition.
Proposition 2.5.
Suppose that $(X,g)$ is a spin Riemannian manifold and $D$ is the associated Dirac operator acting on the spinor bundle. If $\textup{Ind}(D)\neq 0$ in $K_{*}(C^{*}(X)^{\Gamma})$, then zero is in the spectrum of $D$.
For example, the condition $\textup{Ind}(D)\neq 0$ holds when $X$ is geometrically contractible and has finite asymptotic dimension [Yu], or coarsely embedds into Hilbert space [Yucoarseembed].
2.3. Strong Novikov conjecture and its coarse analogue
In this subsection, we recall the statement of the Strong Novikov conjecture for groups and its coarse analogue for non-compact metric spaces.
Let $(X,d)$ be a discrete metric space with bounded geometry. For each $d>0$, we define the Rips complex $P_{d}(X)$ to be the simplical complex generated by points in $X$ such that $x_{i},x_{j}\in X$ are in the same simplex if $d(x_{i},x_{j})\leq d$. By construction, $P_{d}(X)$ is finite dimensional. We equip $P_{d}(X)$ with the spherical metric: for each simplex
$\{\sum_{k=1}^{m}t_{k}x_{i_{k}}:\sum_{k=1}^{n}t_{k}=1,~{}t_{k}\geq 0\},$
its metric is the one obtained from the sphere $\mathbb{S}^{m}$ through the following map:
$\sum_{k=1}^{m}t_{k}x_{i_{k}}\mapsto\Big{(}\frac{t_{0}}{\sqrt{\sum_{k=1}^{n}t_{k}^{2}}},\cdots,\frac{t_{0}}{\sqrt{\sum_{k=1}^{n}t_{k}^{2}}}\Big{)}.$
In particular, if $X=\Gamma$ is a finitely presented group, then $P_{d}(\Gamma)$ admits a natural $\Gamma$-action, which is proper and cocompact. We similarly define the Roe algebra and localization algebra. In particular, the Roe algebras (or the equivariant version) of $X$ and $P_{d}(X)$ are isomorphic [Yulocalization].
Conjecture 2.6 (coarse Novikov conjecture).
Let $X$ be a discrete metric space with bounded geometry. The coarse Novikov conjecture for $X$ states that the evaluation map
$ev\colon\lim_{d\to\infty}C^{*}_{L}(P_{d}(X))\to C^{*}(X)$
induces an injection
$ev_{*}\colon\lim_{d\to\infty}K_{*}(C^{*}_{L}(P_{d}(X)))\to K_{*}(C^{*}(X)).$
Conjecture 2.7 (strong Novikov conjecture).
Let $\Gamma$ be a finitely presented group. The strong Novikov conjecture for $\Gamma$ states that the evaluation map
$ev\colon\lim_{d\to\infty}C^{*}_{L}(P_{d}(\Gamma)^{\Gamma})\to C^{*}_{r}(\Gamma)$
induces an injection
$ev_{*}\colon\lim_{d\to\infty}K_{*}(C^{*}_{L}(P_{d}(\Gamma))^{\Gamma})\to K_{*}(C^{*}_{r}(\Gamma)).$
We briefly list some known results for the above conjectures. The strong Novikov conjecture holds for groups belonging to one of the following cases:
  1. (1)
    groups acting acting properly and isometrically on simply connected and non-positively curved manifolds [GK88],
  2. (2)
    hyperbolic groups [ConnesMoscovici],
  3. (3)
    groups acting properly and isometrically on Hilbert spaces [MR1821144], for example, amenable groups,
  4. (4)
    groups acting properly and isometrically on bolic spaces [MR1998480],
  5. (5)
    groups with finite asymptotic dimension [Yu],
  6. (6)
    groups that coarsely embedds into Hilbert spaces [Yucoarseembed].
The coarse Novikov conjecture holds for metric spaces belonging to one of the following cases:
  1. (1)
    metric spaces that are coarsely equivalent to non-positively curved manifolds [Yulocalization],
  2. (2)
    metric spaces that has finite asymptotic dimension [Yu],
  3. (3)
    metric spaces that coarsely embedds into Hilbert spaces [Yucoarseembed],
In particular, we remark that by the decent principle [Roe], the isomorphism of the map $ev_{*}$ in Conjecture 2.6 for a group $\Gamma$ (as a metric space) implies the strong Novikov conjecture of the group $\Gamma$.
3. Kato inequality for harmonic spinor
The Kato inequality for harmonic spinors is essential for us to obtain the sharpness of the bottom spectrum. In this subsection, we will give a detailed proof inspired by [Davaux, Section 4.1] and [CGH00], in order to give a slightly general version (Proposition 3.3) for non-compact manifolds. For notation simplicity, we only give the proof for real spinors, while the complex case also holds with the same argument.
Proposition 3.1.
Let $(X^{n},g)$ be a complete Riemannian manifold. Let $E$ be a vector bundle over $X$ equipped with a Clifford action of $TX$. Let $D$ be the Dirac operator
$D=\sum_{i=1}^{n}c(e_{i})\nabla_{e_{i}},$
where $\nabla$ is a connection on $E$. If $\xi$ is a smooth section of $E$ such that $D\xi=0$, then
$\Big{|}\nabla|\xi|\Big{|}^{2}\leq\frac{n-1}{n}|\nabla\xi|^{2}.$
Proof.
Let $\nabla\xi$ be the derivative of $\xi$ as a section in $TX\otimes E$. Note that
$\Big{|}\nabla|\xi|^{2}\Big{|}=2\Big{|}\nabla|\xi|\Big{|}\cdot|\xi|=2\Big{|}\langle\nabla\xi,\xi\rangle\Big{|}.$
Therefore, if $\xi(x)\neq 0$ for $x\in X$, then the desired inequality at $x$ is equivalent to
$\Big{|}\langle\nabla\xi(x),\xi(x)\rangle\Big{|}^{2}\leq\frac{n-1}{n}|\nabla\xi(x)|^{2}|\xi(x)|^{2}.$
Since $D\xi(x)=0$, we have $\nabla\xi(x)\in\ker T$, where $T$ is the endomorphism
$T\colon(TX\otimes E)_{x}\to E_{x},~{}\psi\mapsto\sum_{i=1}^{n}c(e_{i})\langle\psi,e_{i}\rangle.$
Now the inequality follows from Lemma 3.2, which will be proved later. Therefore, we have shown that
$\Big{|}\big{(}\nabla|\xi|\big{)}(x)\Big{|}^{2}\leq\frac{n-1}{n}|\big{(}\nabla\xi\big{)}(x)|^{2}$
for any $x\in\mathrm{supp}(\xi)$, namely the support of $\{x\in X:\xi(x)\neq 0\}$. The inequality holds trivially outside $\mathrm{supp}(\xi)$. This finishes the proof. ∎
Lemma 3.2.
Let $V$ be a vector space. Let $W$ be a vector space equipped with a $Cl(V)$-action. Set
$T\colon V\otimes W\to W,~{}\psi\mapsto\sum_{i=1}^{n}c(e_{i})\langle\psi,e_{i}\rangle.$
Then for any $\psi\in\ker T$ and $\xi\in V$, we have
$\Big{|}\langle\psi,\xi\rangle\Big{|}^{2}\leq\frac{n-1}{n}|\psi|^{2}|\xi|^{2}.$
Proof.
Let $\psi=\sum_{i=1}^{n}e_{i}\otimes s_{i}$. Since $\psi\in\ker T$, we have
$\sum_{i=1}^{n}c(e_{i})s_{i}=0.$
Now it suffices to prove that
$\sum_{i=1}^{n}\langle s_{i},\xi\rangle^{2}\leq\frac{n-1}{n}\sum_{i=1}^{n}|s_{i}|^{2}|\xi|^{2}$ (3.1)
subject to the equality for $s_{i}$’s above.
Assume that $|\xi|=1$. We will prove by induction on $n$. The case when $n=1$ is obvious, as $Ts=0$ implies that $s=0$. When $n=2$, we have
$c(e_{1})s_{1}+c(e_{2})s_{2}=0,$
namely $s_{1}=\omega s_{2}$, where $\omega=c(e_{1})c(e_{2})$. Observe that $\omega^{*}=-\omega$ and $\omega^{2}=-1$. Hence
$|s_{2}|=|\omega s_{2}|\text{ and }s_{2}\perp\omega s_{2}.$ (3.2)
It follows that
$\sum_{i=1}^{n}\langle s_{i},\xi\rangle^{2}=\langle\omega s_{2},\xi\rangle^{2}+\langle s_{2},\xi\rangle^{2}=|s_{2}|^{2}\Big{|}P(\xi)\Big{|}^{2}\leq\frac{1}{2}\Big{(}|s_{1}|^{2}+|s_{2}|^{2}\Big{)}|\xi|^{2},$
where $P$ is the orthogonal projection from $W$ to $\text{span}\{s_{2},\omega s_{2}\}$. This finishes the proof when $n=2$. In particular, the equality holds if and only if $\text{span}\{s_{2},\omega s_{2}\}$ or $s_{2}=0$.
Now we prove the inequality (3.1) for $n\geq 3$ by induction. For any $i\neq j$, we define
$s_{i,j}=s_{i}-\frac{1}{n-1}c(e_{i})c(e_{j})s_{j}.$
Since $\sum_{i=1}^{n}c(e_{i})s_{i}=0$, we see that $\sum_{i:i\neq j}c(e_{i})s_{i,j}=0.$ By the induction hypothesis, we have
$\sum_{i:i\neq j}\langle s_{i,j},\xi\rangle^{2}\leq\frac{n-2}{n-1}\sum_{i:i\neq j}|s_{i,j}|^{2}.$
Take summation for $j=1,2,\ldots,n$ and obtain that
$\sum_{i\neq j}\langle s_{i,j},\xi\rangle^{2}\leq\frac{n-2}{n-1}\sum_{i\neq j}|s_{i,j}|^{2}$ (3.3)
We first compute the right-hand side of line (3.3). Note that
$|s_{i,j}|^{2}=$ $\Big{|}s_{i}-\frac{1}{n-1}c(e_{i})c(e_{j})s_{j}\Big{|}^{2}=|s_{i}|^{2}+\frac{1}{(n-1)^{2}}|s_{j}|^{2}-\frac{2}{n-1}\langle s_{i},c(e_{i})c(e_{j})s_{j}\rangle$
$=$ $|s_{i}|^{2}+\frac{1}{(n-1)^{2}}|s_{j}|^{2}+\frac{2}{n-1}\langle c(e_{i})s_{i},c(e_{j})s_{j}\rangle$
Sum for $i$ with $i\neq j$,
$\sum_{i:i\neq j}|s_{i,j}|^{2}=\sum_{i:i\neq j}|s_{i}|^{2}+\frac{1}{n-1}|s_{j}|^{2}-\frac{2}{n-1}|s_{j}|^{2}=\sum_{i:i\neq j}|s_{i}|^{2}-\frac{1}{n-1}|s_{j}|^{2}.$
It follows that
$\begin{split}\sum_{i\neq j}|s_{i,j}|^{2}=&\sum_{i\neq j}|s_{i}|^{2}-\frac{1}{n-1}\sum_{j=1}^{n}|s_{j}|^{2}=(n-1)\sum_{i=1}^{n}|s_{i}|^{2}-\frac{1}{n-1}\sum_{j=1}^{n}|s_{j}|^{2}\\ =&\frac{n(n-2)}{n-1}\sum_{i=1}^{n}|s_{i}|^{2}.\end{split}$ (3.4)
Now we estimate the left-hand side of line (3.3). Fix $i\in\{1,\ldots,n\}$. We have
$\sum_{j:i\neq j}\langle s_{i,j},\xi\rangle^{2}=\sum_{j:i\neq j}\langle s_{i}-\frac{1}{n-1}c(e_{i})c(e_{j})s_{j},\xi\rangle^{2}$
By the Cauchy–Schwarz inequality, we have
$\begin{split}&(n-1)\sum_{j:i\neq j}\langle s_{i}-\frac{1}{n-1}c(e_{i})c(e_{j})s_{j},\xi\rangle^{2}\\ \geq&\Big{(}\sum_{j:i\neq j}\langle s_{i}-\frac{1}{n-1}c(e_{i})c(e_{j})s_{j},\xi\rangle\Big{)}^{2}\\ =&\Big{\langle}(n-1)s_{i}-\frac{1}{n-1}\sum_{j:i\neq j}c(e_{i})c(e_{j})s_{j},\xi\Big{\rangle}^{2}\\ =&\frac{n^{2}(n-2)^{2}}{(n-1)^{2}}\langle s_{i},\xi\rangle^{2}.\end{split}$ (3.5)
Here the last equality follows from
$\sum_{j:i\neq j}c(e_{i})c(e_{j})s_{j}=c(e_{i})\Big{(}-c(e_{i})s_{i}\Big{)}s_{i}=s_{i}.$
Thus we obtain an estimate for the left-hand side of line (3.3) as
$\sum_{i\neq j}\langle s_{i,j},\xi\rangle^{2}\geq\frac{n^{2}(n-2)^{2}}{(n-1)^{3}}\sum_{i=1}^{n}\langle s_{i},\xi\rangle^{2}.$ (3.6)
Combining (3.3), (3.4), and (3.6), we obtain that
$\frac{n^{2}(n-2)^{2}}{(n-1)^{3}}\sum_{i=1}^{n}\langle s_{i},\xi\rangle^{2}\leq\frac{n-2}{n-1}\cdot\frac{n(n-2)}{n-1}\sum_{i=1}^{n}|s_{i}|^{2}.$
Since $n\geq 3$, we have
$\sum_{i=1}^{n}\langle s_{i},\xi\rangle^{2}\leq\frac{n-1}{n}\sum_{i=1}^{n}|s_{i}|^{2}.$
This finishes the proof. ∎
Indeed, the proof of Lemma 3.2 implies a slightly general version of Proposition 3.1 as follows.
Proposition 3.3.
Let $E$ be a bundle over $X^{n}$ equipped with a Clifford action of $TX$. There exists $c_{n}>0$ that only depends on $n$ such that for any smooth section $\xi$ of $E$, we have
$\Big{|}\nabla|\xi|\Big{|}^{2}\leq\frac{n-1}{n}|\nabla\xi|^{2}+c_{n}|D\xi|^{2}+c_{n}|D\xi||\nabla\xi|.$
4. Sharp Bottom spectrum and scalar curvature rigidity
In this section, we will prove the main theorems and then state several related corollaries. Let us first recall the definition of a net in a metric space.
Definition 4.1.
Suppose that $(X,d)$ is a metric space and $S$ is a subset of $X$. $S$ is said to be a net of $X$ if there exists $r>0$ such that $N_{r}(S)=X$, where $N_{r}(S)=\{x\in X:\mathrm{dist}(x,S)<r\}$. Furthermore, we say that $S$ is a discrete net of $X$ if there exists $r^{\prime}>0$ such that $d(x,y)\geq r^{\prime}$ for any $x\neq y$ in $S$.
Theorem 4.2.
Suppose that $(X^{n},g)$ is a complete, noncompact, spin Riemannian manifold with bounded geometry and $D$ is the Dirac operator acting on the spinor bundle over $X$. If
  1. (1)
    $\textup{Ind}(D)\in K_{*}(C^{*}(X))$ is non-zero, and
  2. (2)
    $\mathrm{Sc}_{g}\geq-\kappa$ for some constant $\kappa\geq 0$,
then
$\lambda_{1}(X,g)\leq\frac{n-1}{4n}\kappa.$
Moreover, if $\lambda_{1}(X,g)=\frac{n-1}{4n}\kappa$, then for any $\delta>0$, the set
$\{x\in X:\mathrm{Sc}_{g}(x)\geq-\kappa+\delta\}$
is not a net of $X$.
Theorem 4.3.
Suppose that $(M^{n},g)$ is a closed spin Riemannian manifold and $(\widetilde{M},\widetilde{g})$ is the Riemannian universal cover of (M, g). Let $\Gamma=\pi_{1}(M)$ and $\widetilde{D}$ be the Dirac operator acting on the spinor bundle over $\widetilde{M}$. If
  1. (1)
    $\textup{Ind}(\widetilde{D})\in K_{*}(C^{*}(\widetilde{M})^{\Gamma})\cong K_{*}(C^{*}_{r}(\Gamma))$ is non-zero;
  2. (2)
    $\mathrm{Sc}_{g}\geq-\kappa$ for some constant $\kappa\geq 0$,
then
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{n-1}{4n}\kappa.$
Moreover, if $\lambda_{1}(\widetilde{M},\widetilde{g})=\frac{n-1}{4n}\kappa$, then $(M^{n},g)$ has constant scalar curvature $\mathrm{Sc}_{g}=-\kappa$ on $M$.
The proof of sharpness is classical, originating from the methods in [Davaux], where the technique is integrated with classical index theory. However, the net characterization for complete noncompact manifolds and the scalar curvature rigidity characterization for cocompact manifolds in Theorem 1.1 and Theorem 1.3 are innovative. To our best knowledge, they have not previously appeared in the geometric analysis literature under the assumption of scalar curvature lower bound. The key element for the rigidity results is the following unique continuation inequality, which is quite technical. Therefore, the proof is deferred to Section 5. Here, we present only the version involving the Dirac operator, which is necessary for our purposes in this section. For a more general case and its proof, see Section 5.
Proposition 4.4.
Suppose that $(X^{n},g)$ is a complete spin Riemannian manifold with bounded geometry. Let $Y$ be a discrete net of $X$ and $N_{a}(Y)$ the $a$-neighborhood of $Y$ for some $a>0$, $S$ be the spinor bundle over $X$, and $D$ the Dirac operator. If $P_{\lambda}$ is the spectral projection of $P$ acting on $L^{2}(S)$ with spectrum $\leq\lambda$ and $V_{\lambda}$ is the range of $P_{\lambda}$, then there exists a constant $C_{\lambda}>0$ such that
$\int_{X}|\sigma|^{2}\leq C_{\lambda}\int_{N_{a}(Y)}|\sigma|^{2}\text{ for any }\sigma\in V_{\lambda}.$
Proof of Theorem 4.2.
Let $S_{X}$ be the spinor bundle over $X$. Since $\textup{Ind}(D)\in K_{*}(C^{*}(X))$ is non-zero, we obtain that the Dirac operator $D$ is not invertible. Consequently, for any $\varepsilon>0$, there exists a spinor $s\in L^{2}(S_{X})$ such that
$\|s\|=1\text{ and }\|Ds\|\leq\varepsilon.$
Note that
  • The Lichnerowicz formula shows that
    $\|\nabla s\|^{2}=\|Ds\|^{2}-\int_{X}\frac{\mathrm{Sc}_{g}}{4}|s|^{2}\leq\varepsilon^{2}+\frac{\kappa}{4}.$
  • The Kato inequality in Proposition 3.3 indicates that there exists $c_{n}>0$ such that
    $\Big{|}\nabla|s|\Big{|}^{2}\leq\frac{n-1}{n}|\nabla s|^{2}+c_{n}|Ds|^{2}+c_{n}|Ds||\nabla s|$
    in $(X,{g})$.
By Integrating on $X$, we obtain that
$\int_{X}\langle-\Delta|s|,|s|\rangle=\Big{\|}\nabla|s|\Big{\|}^{2}\leq\frac{n-1}{n}\big{(}\varepsilon^{2}+\frac{\kappa}{4}\big{)}+c_{n}\varepsilon^{2}+c_{n}\varepsilon\sqrt{\varepsilon^{2}+\frac{\kappa}{4}}.$
Since $\varepsilon$ can be picked as any positive real number, we let $\varepsilon\rightarrow 0$ and then we obtain
$\lambda_{1}(X,g)\leq\frac{n-1}{4n}\kappa.$
Next, let us prove the scalar curvature rigidity if the equality holds as follows. We will argue by contradication. Suppose that there exists a positive constant $\delta>0$ such that the set
$X_{\delta}:=\{x\in X:\mathrm{Sc}_{g}(x)\geq-\kappa+\delta\}$
is a net of $X$, then there exists a discrete net $Y$ of $X$ and some $a>0$ such that
$\mathrm{Sc}_{g}(x)\geq-\kappa+\delta\text{ for any }x\in N_{a}(Y).$
Here, we have used the assumption of bounded geometry.
Now given any $\varepsilon>0$, let $P_{\varepsilon^{2}}$ be the spectral projection to the spectrum $\leq\varepsilon^{2}$ and $V_{\varepsilon^{2}}$ the range of $P_{\varepsilon^{2}}$. Since $D$ is non invertible, we obtain that $V_{\varepsilon^{2}}$ is non-empty. Let us pick a spinor $s$ in $V_{\varepsilon^{2}}$ with $\|s\|=1$. Clearly we have $\|Ds\|\leq\varepsilon.$ By our assumption that $\mathrm{Sc}_{g}\geq-\kappa+\delta$ on $N_{a}(Y)$, we obtain by the Lichnerowicz formula that
$\|\nabla s\|^{2}=\|Ds\|^{2}-\int_{X}\frac{\mathrm{Sc}_{g}}{4}|s|^{2}\leq\varepsilon^{2}+\frac{\kappa}{4}-\frac{\delta}{4}\|s\|^{2}_{L^{2}(N_{a}(Y))}.$
Similarly, we deduce
$\Big{\|}\nabla|s|\Big{\|}^{2}\leq\frac{n-1}{n}\big{(}\varepsilon^{2}+\frac{\kappa}{4}\big{)}+c_{n}\varepsilon^{2}+c_{n}\varepsilon\sqrt{\varepsilon^{2}+\frac{\kappa}{4}}-\frac{(n-1)\delta}{4n}\|s\|^{2}_{L^{2}(N_{a}(Y))}.$
Assume that $\varepsilon<1$. By Proposition 4.4, there exists $C>0$ independent of $\varepsilon$ such that
$\|s\|_{L^{2}(N_{a}(Y))}\geq\frac{1}{C}\|s\|=\frac{1}{C}.$
Therefore, we see that
$\Big{\|}\nabla|s|\Big{\|}^{2}\leq\frac{(n-1)\kappa}{4n}-\frac{(n-1)\delta}{4nC^{2}}+\Big{(}\frac{n-1}{n}\varepsilon^{2}+c_{n}\varepsilon^{2}+c_{n}\varepsilon\sqrt{\varepsilon^{2}+\frac{\kappa}{4}}\Big{)}.$
By letting $\varepsilon\to 0$, we have
$\lambda_{1}(X,g)\leq\frac{n-1}{4n}\kappa-\frac{(n-1)\delta}{4nC^{2}}<\frac{n-1}{4n}\kappa.$
This contradicts with the assumption that $\lambda_{1}(X,g)=\frac{n-1}{4n}\kappa$. This finishes the proof. ∎
We also list some topological conditions where $\textup{Ind}(D)\in K_{*}(C^{*}(X))$ is non-zero.
Proposition 4.5.
If $(X,g)$ is a geometrically contractible Riemannian manifold and satisfies the coarse Novikov conjecture , then $\textup{Ind}(D)$ in non-zero in $K_{*}(C^{*}(X))$. In particular, Theorem 1.1 holds for $X$.
In particular, the above property holds for asymptotically hyperbolic manifolds. A complete Riemannian manifold $(X^{2},g)$ is said to be asymptotically hyperbolic if it is conformally compact with the standard sphere $(\mathbb{S}^{n-1},g_{\mathbb{S}^{n-1}})$ as its conformal end, and there is a unique defining function $r$ in a collar neighborhood near infinite such that
$g=\sinh^{-2}(r)\big{(}dr^{2}+g_{\mathbb{S}^{n-1}}+\frac{r^{n}}{n}h+O(r^{n+1})\big{)}$
where $h$ is a symmetric 2-tensor on $\mathbb{S}^{n-1}$, and the asymptotic expression can be differentiated twice. See [MR1879228]. Therefore, we have the following corollary.
Corollary 4.6.
Suppose that $(X^{n},g)$ is an asymptotically hyperbolic and spin manifold with scalar curvature $\mathrm{Sc}_{g}\geq-n(n-1)$, then
$\lambda_{1}({X}^{n},g)\leq\frac{(n-1)^{2}}{4}.$
Moreover, the coarse Novikov conjecture holds for Riemannian manifold with non-positive sectional curvature (see [Yulocalization]). Hence, a geometric version of Theorem 4.2 is as follows.
Corollary 4.7.
Suppose that $(X^{n},g)$ is a complete manifold with non-positive sectional curvature $\sec_{g}\leq 0$ and bounded geometry. If $\mathrm{Sc}_{g}\geq-n(n-1)\kappa$, then
$\lambda_{1}(X,g)\leq\frac{(n-1)^{2}}{4}\kappa.$
Finally, let us prove Theorem 4.3. The proof is similar to that of Theorem 4.2 so we omit most of the details.
Proof of Theorem 4.3.
Let $\widetilde{S}$ be the spinor bundle over $(\widetilde{M},\widetilde{g})$ and $\widetilde{D}$ the Dirac operator acting on $\widetilde{S}$. If $\textup{Ind}(\widetilde{D})$ is non-zero in $K_{*}(C^{*}_{r}(\Gamma))$, then the operator $\widetilde{D}$ is not invertible. Therefore, by the same argument using the Lichnerowicz formula and the Kato inequality, we obtain that
$\lambda_{1}(\widetilde{M},\widetilde{g})\leq\frac{n-1}{4n}\kappa.$
Furthermore, let us assume that $\lambda_{1}(\widetilde{M},\widetilde{g})=\frac{n-1}{4n}\kappa$ and $\mathrm{Sc}_{g}$ is not equal to $\kappa$ everywhere on $M$. More precisely, assume that there exists $x_{0}\in M$ , $\delta>0$, and $a>0$, such that $\mathrm{Sc}_{g}\geq\kappa+\delta$ on $N_{a}(x_{0})\subset M$. Let $Y$ be the lift of the point $\{x_{0}\}$ in $\widetilde{M}$, which is a discrete net of $\widetilde{M}$. Therefore, as $\widetilde{g}$ is a lifted metric, $\mathrm{Sc}_{\widetilde{g}}\geq\kappa+\delta$ on $N_{a}(Y)$. This contradicts with the assumption that $\lambda_{1}(\widetilde{M},\widetilde{g})=\frac{n-1}{4n}\kappa$ by the same proof of Theorem 4.2 using Proposition 4.4. ∎
We remark that the index theoretical condition, namely $\textup{Ind}(D)\in K_{*}(C^{*}(X)^{\Gamma})$ is non-zero, can be verified under the following topological and algebraic condition:
Proposition 4.8.
Suppose that $(M^{n},g)$ is a closed spin Riemannian manifold and $\Gamma=\pi_{1}(M)$. If
  • $X$ is rationally essential, namely the fundamental class $[M]$ is non-zero in $H_{*}(B\Gamma,\mathbb{Q})$, and
  • $\Gamma$ satisfies the strong Novikov conjecture 2.7,
then $\textup{Ind}(\widetilde{D})$ in non-zero in $K_{*}(C^{*}_{r}(\Gamma))$. Hence, Theorem 1.3 holds.
5. Unique continuation Theorem on complete Riemannian manifold
Suppose that $(X^{n},g)$ is a complete Riemannian manifold with bounded geometry and $P$ is a second order elliptic differential operator on $X$ acting on a smooth bundle $E$ over $X$. Then, the elliptic operator theory shows that $P$ satisfies the Gårding’s inequality. Namely, there exists constants $c,c^{\prime}>0$ such that
$\langle P\sigma,\sigma\rangle\geq c\|\nabla\sigma\|^{2}-c^{\prime}\|\sigma\|^{2}.$ (5.1)
Let $Y$ be a discrete subset of $X$. In this section, we will prove the following unique continuation theorem.
Theorem 5.1.
Let $(X^{n},g)$ be a complete Riemannian manifold with bounded geometry. Let $Y$ be a discrete net of $X$ and $N_{a}(Y)$ the $a$-neighborhood of $Y$ for some $a>0$. Let $E$ be a vector bundle over $X$ and $P$ a second order elliptic differential operator acting on $E$ satisfying the Gårding inequality in line (5.1). If $P_{\lambda}$ is the spectral projection of $P$ acting on $L^{2}(E)$ with spectrum $\leq\lambda$ and $V_{\lambda}$ is the range of $P_{\lambda}$, then there exists a constant $C_{\lambda}>0$ such that
$\|\sigma\|_{L^{2}(X)}\leq C_{\lambda}\|\sigma\|_{L^{2}(N_{a}(Y))}\text{ for any }\sigma\in V_{\lambda}.$
Theorem 5.1 is essentially motivated by [MR1362555, MR3041662]. It plays an inessential role in the proof of main theorem regarding the scalar curvature rigidity/scalar curvature distribution.
5.1. Local Carleman estimate
In this subsection, we will prove a local Carleman estimate for elliptic differential operators on a discrete net in complete manifold $(X^{n},g)$.
Let $X\times\mathbb{R}_{\geq 0}$ be the product space of $X$ and the half real line. In the following proof, we will use function $\varphi$ as key ingredients in variant circumstances. To begin with, we consider the simple case when $y$ is a singleton in $X$ and give a detailed computation. Given any fixed point $y\in X$, we consider a function on $X\times\mathbb{R}_{\geq 0}$
$\varphi(x,t)=e^{-t-d(x,y)^{6}},$
whose derivatives along $X$ with order $\leq 5$ are small near $y$.
Given a fixed small $a>0$, let $\mathcal{F}$ be the space of smooth sections in $E$ over $X\times\mathbb{R}_{\geq 0}$ that are supported in $\{\varphi<a\}$ and vanish on $X\times\{0\}$. Let
$Q=-\frac{\partial^{2}}{\partial t^{2}}+P$
be a differential operator that acts on $\mathcal{F}$. For any $h>0$, we define
$Q_{\varphi}=e^{\varphi/h}\cdot Q\cdot e^{-\varphi/h}.$
We first prove
Lemma 5.2.
There exists $C_{1},C_{2}>0$ such that for any $f\in\mathcal{F}$, we have
$\frac{1}{h}\|\frac{\partial f}{\partial t}\|^{2}+\frac{1}{h}\|\nabla f\|^{2}+\frac{1}{h^{3}}\|f\|^{2}\leq C_{1}\|Q_{\varphi}f\|^{2}+\frac{C_{2}}{h}\int_{X\times\{0\}}|\frac{\partial f}{\partial t}|^{2}$
for any $h>0$ sufficiently small.
Proof.
Let $A$ and $B$ be the self-adjoint and anti-self-adjoint parts of $Q_{\varphi}$ respectively, namely
$A=\frac{Q_{\varphi}+Q_{\varphi}^{*}}{2},\ B=\frac{Q_{\varphi}-Q_{\varphi}^{*}}{2}.$
A direct calculation shows that
$A=Q-\frac{\dot{\varphi}^{2}}{h^{2}}+\mathscr{R}_{1}=-\frac{\partial^{2}}{\partial t^{2}}+P-\frac{\dot{\varphi}^{2}}{h^{2}}+\mathscr{R}_{1},$
$B=2\frac{\dot{\varphi}}{h}\frac{\partial}{\partial t}+\frac{\ddot{\varphi}}{h}+\mathscr{R}_{2}=\frac{\partial}{\partial t}\frac{\dot{\varphi}}{h}+\frac{\dot{\varphi}}{h}\frac{\partial}{\partial t}+\mathscr{R}_{2}.$
Here, we denote by $\dot{\varphi},\ddot{\varphi}$ the derivatives of $\varphi$ with respect to $t\in\mathbb{R}$, and $\mathscr{R}_{1}$ and $\mathscr{R}_{2}$ are the remainders given by the derivatives of $\varphi$ along $X$, which are small by the construction.
Note that $Q_{\varphi}=A+B$, we have
$\|Q_{\varphi}f\|^{2}=\|Af\|^{2}+\|Bf\|^{2}+\langle Af,Bf\rangle+\langle Bf,Af\rangle.$
Since $f$ is compactly supported within $X\times[0,a)$ and $f(x,0)=0$ for any $x\in X$, we have
$\langle Af,Bf\rangle=-\langle BAf,f\rangle$
and
$\langle Bf,Af\rangle=\langle ABf,f\rangle-\langle Bf,\frac{\partial f}{\partial t}\rangle\Big{|}_{0}^{a}=\langle ABf,f\rangle+\int_{X\times\{0\}}\frac{2\dot{\varphi}}{h}|\frac{\partial f}{\partial t}|^{2}.$
It follows that
$\|Q_{\varphi}f\|^{2}-\int_{X\times\{0\}}\frac{2\dot{\varphi}}{h}|\frac{\partial}{\partial t}f|^{2}=\|Af\|^{2}+\|Bf\|^{2}+\langle[A,B]f,f\rangle.$
Here $[A,B]=AB-BA$. A direct computation shows that
$[A,B]=$ $[-\frac{\partial^{2}}{\partial t^{2}}-\frac{\dot{\varphi}^{2}}{h^{2}},\frac{\partial}{\partial t}\frac{\dot{\varphi}}{h}+\frac{\partial}{\partial t}\frac{\dot{\varphi}}{h}]+\mathscr{R}_{3}$
$=$ $4\frac{\dot{\varphi}^{2}\ddot{\varphi}}{h^{3}}-\frac{\partial}{\partial t}\Big{(}2\frac{\ddot{\varphi}}{h}\frac{\partial}{\partial t}+\frac{\dddot{\varphi}}{h}\Big{)}-\Big{(}2\frac{\ddot{\varphi}}{h}\frac{\partial}{\partial t}+\frac{\dddot{\varphi}}{h}\Big{)}\frac{\partial}{\partial t}+\mathscr{R}_{3}.$
Here the remainder $\mathscr{R}_{3}$ is also small and will be ignored. By construction, we have $\dot{\varphi}^{2}\ddot{\varphi}>1/2$ and $\ddot{\varphi}>1/2$ on the support of $f$ if $a$ is small enough. Furthermore, by line (5.1) and Cauchy–Schwarz inequality, we have
$\langle\frac{\dot{\varphi}^{2}}{h^{3}}f,f\rangle=$ $\frac{1}{h}\langle-\frac{\partial^{2}}{\partial t^{2}}f+Pf-Af+\mathscr{R}_{1}f,f\rangle$
$\geq$ $\frac{1}{h}\|\frac{\partial}{\partial t}f\|^{2}+\frac{1}{h}c\|\nabla f\|^{2}-\frac{1}{h}c^{\prime}\|f\|^{2}$
$-\frac{h^{1/2}}{2}\|Af\|^{2}-\frac{1}{2h^{5/2}}\|f\|^{2}+\frac{1}{h}\langle\mathscr{R}_{1}f,f\rangle,$
and
$\langle-\frac{\partial}{\partial t}\Big{(}2\frac{\ddot{\varphi}}{h}\frac{\partial}{\partial t}+\frac{\dddot{\varphi}}{h}\Big{)}f-\Big{(}2\frac{\ddot{\varphi}}{h}\frac{\partial}{\partial t}+\frac{\dddot{\varphi}}{h}\Big{)}\frac{\partial}{\partial t}f,f\rangle$
$=$ $\frac{1}{h}\langle{4\ddot{\varphi}}\frac{\partial}{\partial t}f,\frac{\partial}{\partial t}f\rangle+\frac{1}{h}\langle 2\dddot{\varphi}f,\frac{\partial}{\partial t}f\rangle$
$\geq$ $\frac{2}{h}\|\frac{\partial}{\partial t}f\|^{2}-\frac{1}{h^{1/2}}\|\frac{\partial}{\partial t}f\|^{2}-\frac{1}{h^{3/2}}\langle|\dddot{\varphi}|^{2}f,f\rangle.$
Hence there exists $c_{1}>0$ such that
$\langle[A,B]f,f\rangle\geq c_{1}\Big{(}\frac{1}{h}\|\frac{\partial}{\partial t}f\|^{2}+\frac{1}{h}\|\nabla f\|^{2}+\frac{1}{h^{3}}\|f\|^{2}\Big{)}-c_{1}\sqrt{h}\|Af\|^{2}.$ (5.2)
This inequality indicates that $[A,B]$ is positive modulo $A$. Clearly $\|Bf\|^{2}\geq 0$. This finishes the proof for $h$ sufficiently small. ∎
We remark that the key ingredient that proves Lemma 5.2 is the non-negativity condition (5.2). This follows from the fact that the function $e^{x}$ has positive second order derivative, and the function $\varphi$ has non-zero derivative along some direction. The estimate (5.2) holds more generally if $\varphi$ satisfies Hörmander’s condition [HormanderIV, Theorem 27.1.11]. See [MR3041662, Section 3]. The following lemma is directly from Lemma 5.2 by substituting $f=e^{\varphi/h}g$.
Lemma 5.3.
There exists $C_{1},C_{2}>0$ such that for any $g\in\mathcal{F}$, we have
$\int_{X\times\mathbb{R}_{\geq 0}}\Big{(}\frac{1}{h}|\frac{\partial g}{\partial t}|^{2}+\frac{1}{h}|\nabla g|^{2}+\frac{1}{h^{3}}|g|^{2}\Big{)}e^{2\varphi/h}$
$\leq$ $C_{1}\int_{X\times\mathbb{R}_{\geq 0}}e^{2\varphi/h}|Qg|^{2}+\frac{C_{2}}{h}\int_{X\times\{0\}}|\frac{\partial g}{\partial t}|^{2}.$
for any $h>0$ sufficiently small.
Now we consider the non-compact case. Let $Y$ be a discrete net of $X$. Given any fixed small $a>0$, let $\mathcal{F}_{Y}$ be the space of smooth sections $g$ of $E$ over $X\times\mathbb{R}_{\geq 0}$ that satisfy
  • $g$ is supported in $N_{a}(Y)\times[0,a)$,
  • $g|_{X\times\{0\}}=0$.
Let $\varphi_{Y}$ be a function on $X\times\mathbb{R}_{\geq 0}$ defined by
$\varphi_{Y}(x,t)=e^{-t-d(x,y)^{6}}$
on $B_{2a}(y)\times[0,2a]$ for any $y\in Y$. We assume that $a$ is small enough so that the $4a$-neighborhoods of points in $Y$ are disjoint in Definition 4.1. The value of $\varphi_{Y}$ on the rest of points in $X$ can be arbitrary.
As it is pointed out in line (5.2), the essential part for the proof of Lemma 5.2 is the non-negativity condition (5.2) on the support of $f$. Here, we note that if $f\in\mathcal{F}_{Y}$, line (5.2) holds for $\varphi_{Y}$ as well. Thus, Lemma 5.3 still holds for the non-compact case.
Lemma 5.4.
There exists $C_{1},C_{2}>0$ such that for any $g\in\mathcal{F}_{Y}$, we have
$\begin{split}&\int_{X\times\mathbb{R}_{\geq 0}}\Big{(}\frac{1}{h}|\frac{\partial g}{\partial t}|^{2}+\frac{1}{h}|\nabla g|^{2}+\frac{1}{h^{3}}|g|^{2}\Big{)}e^{2\varphi_{Y}/h}\\ \leq&C_{1}\int_{X\times\mathbb{R}_{\geq 0}}e^{2\varphi_{Y}/h}|Qg|^{2}+\frac{C_{2}}{h}\int_{X\times\{0\}}|\frac{\partial g}{\partial t}|^{2}\\ \end{split}$ (5.3)
for any $h>0$ sufficiently small.
Moreover, we will consider another type of function $\varphi$ along directions in $X$. Given a discrete net $Y$ of $X$, let $Z=\{Z_{i}\}$ be a collection of pieces of oriented hypersurfaces, where each piece is located near a point of $Y$. We fix a small number $t_{0}>0$, and points $z_{i}\in Z_{i}$. Pick smooth functions $v_{i}$ supported near $Z_{i}$ such that $|\nabla v_{i}|=1$, $Z_{i}$ is the level set $\{v_{i}=0\}$, and $\nabla v_{i}$ is pointing outward from $Z_{i}$. We define $\varphi_{Z}$ on $X\times\mathbb{R}_{\geq 0}$ as
$\varphi_{Z}(x,t)=-v_{i}-d((x,t),(z_{i},t_{0}))^{6}$
near each $Z_{i}$. The value of $\varphi$ away from $Z_{i}$ is arbitrary.
Let $\mathcal{F}_{Z}$ be the collections of smooth sections of $E$ over $X\times\mathbb{R}_{\geq 0}$ that are supported in a small neighborhood of $Z\times\{t_{0}\}$. The same proof of Lemma 5.4 applies to the function $\varphi_{Z}$.
Lemma 5.5.
There exists $C_{1}>0$ such that for any $g\in\mathcal{F}_{Z}$, we have
$\int_{X\times\mathbb{R}_{\geq 0}}\Big{(}\frac{1}{h}|\frac{\partial g}{\partial t}g|^{2}+\frac{1}{h}|\nabla g|^{2}+\frac{1}{h^{3}}|g|^{2}\Big{)}e^{2\varphi_{Z}/h}\leq C_{1}\int_{X\times\mathbb{R}_{\geq 0}}e^{2\varphi_{Z}/h}|Qg|^{2}$ (5.4)
for any $h>0$ sufficiently small.
Note that the calculation in Lemma 5.2 applies to the function $\varphi_{Z}$ if we replace the $t$-direction derivative by the $\nabla v_{i}$-directions. Thus, Lemma 5.5 follows from the similar calculation. We also emphasize that as $g$ is supported near $Z$, we only need the value of the function $\varphi_{Z}$ near $Z$ in the proof. Since $g$ vanishes away from $Z$, the boundary term, namely the second term in the right-hand side of line (5.3), does not appear in line (5.4).
5.2. Interpolation and unique continuation
In this subsection, we will first prove an interpolation inequality for sections over $X\times\mathbb{R}_{\geq 0}$ and then derive the unique continuation theorem at lower spectrum of elliptic operators from a net.
We begin with some elementary inequalities that deduce interpolation inequality from a Carleman estimate.
Lemma 5.6.
Let $\alpha,\beta,\gamma$ be positive numbers with $\alpha\leq A\beta$ for some $A>0$. If there exist $p,q>0$ and $h_{0}>0$ such that
$\alpha\leq e^{-p/h}\beta+e^{q/h}\gamma$
for any $h\in(0,h_{0})$, then there exist $C>0$ and $\nu\in(0,1)$ that only depends on $A,p,q,h_{0}$ such that
$\alpha\leq C\beta^{\nu}\gamma^{1-\nu}.$
Proof.
We set $\nu=\frac{q}{p+q}$ and define the function $F(h)=e^{-p/h}\beta+e^{q/h}\gamma$ on $\mathbb{R}^{+}$. A direct calculation shows that $F$ attains its unique minimum at the point
$h=h_{*}=\frac{\ln(p\beta)-\ln(q\gamma)}{p+q},$
and the minimum value is
$F(h_{*})=(p+q)p^{-\frac{p}{p+q}}q^{-\frac{q}{p+q}}\cdot\beta^{\nu}\gamma^{1-\nu}.$
Let us consider the following cases.
  • If $h_{*}\leq h_{0}$, then the desired inequality follows direclty.
  • Assume that $h_{*}\geq h_{0}$.
    • If $\beta\leq\gamma$, then we have obviously
      $\alpha\leq A\beta\leq A\beta^{\nu}\gamma^{1-\nu}.$
    • If $\gamma\leq\beta$, then by the monotonicity of $F$ on $(0,h_{0})$, we have
      $\alpha$ $\leq F(h_{0})=e^{-p/h_{0}}\beta+e^{q/h_{0}}\gamma$
      $\leq e^{-p/h_{*}}\beta+e^{q/h_{0}}\gamma=q^{\frac{p}{p+q}}p^{-\frac{p}{p+q}}\beta^{\nu}\gamma^{1-\nu}+e^{q/h_{0}}\gamma$
      $\leq\Big{(}q^{\frac{p}{p+q}}p^{-\frac{p}{p+q}}+e^{q/h_{0}}\Big{)}\beta^{\nu}\gamma^{1-\nu},$
      where the last inequality follows from $\gamma\leq\beta$.
To summarize, we have shown that $\alpha\leq C\beta^{\nu}\gamma^{1-\nu}$ by setting
$C=\max\Big{\{}(p+q)p^{-\frac{p}{p+q}}q^{-\frac{q}{p+q}},A,q^{\frac{p}{p+q}}p^{-\frac{p}{p+q}}+e^{q/h_{0}}\Big{\}}.$
Lemma 5.7.
Suppose that $\alpha_{i}>0$ for $i=0,1,\cdots,N$ and $\beta,\gamma$ are positive numbers with $\alpha_{i}\leq\beta$ for any $i$. If there exists $\nu\in(0,1)$ and $C\geq 1$ such that
$\alpha_{k+1}\leq C\beta^{\nu}(\alpha_{k}+\gamma)^{1-\nu},\ k=0,\cdots,N-1$
then,
$\alpha_{N}\leq C^{\prime}\beta^{\mu}(\alpha_{0}+\gamma)^{1-\mu},$
where $\mu=1-(1-\nu)^{N}$ and $C^{\prime}=(2C)^{1+(1-\nu)+\cdots+(1-\nu)^{N-1}}$.
Proof.
If $\gamma\geq\beta$, then $\alpha_{0}+\gamma\geq\beta$, then
$\alpha_{N}\leq\beta\leq\beta^{\mu}(\alpha_{0}+\gamma)^{1-\mu}.$
Now we assume $\gamma\leq\beta$, then we obtain that
$\frac{\gamma}{\beta}\leq\frac{\gamma^{1-\nu}}{\beta^{1-\nu}}\leq\frac{(\alpha_{k}+\gamma)^{1-\nu}}{\beta^{1-\nu}}\leq C\frac{(\alpha_{k}+\gamma)^{1-\nu}}{\beta^{1-\nu}}$
for any $k$. Moreover, the assumption implies that
$\frac{\alpha_{k+1}}{\beta}\leq C\frac{(\alpha_{k}+\gamma)^{1-\nu}}{\beta^{1-\nu}}.$
Hence, we reach
$\frac{\alpha_{k+1}+\gamma}{\beta}\leq 2C\frac{(\alpha_{k}+\gamma)^{1-\nu}}{\beta^{1-\nu}}.$
Therefore,
$\frac{\alpha_{N}}{\beta}\leq 2C\frac{(\alpha_{N-1}+\gamma)^{1-\nu}}{\beta^{1-\nu}}\leq\cdots\leq C^{\prime}\frac{(\alpha_{0}+\gamma)^{(1-\nu)^{N}}}{\beta^{(1-\nu)^{N}}}.$
Equivalently,
$\alpha_{N}\leq C^{\prime}\beta^{\mu}(\alpha_{0}+\gamma)^{1-\mu}.$
This finishes the proof. ∎
We start from the following lemma by applying the construction in Lemma 5.4 first.
Lemma 5.8.
Suppose that $X$ is complete Riemannian manifld, and $Y$ is a discrete net in $X$ and $a$ is a small positive number. Given small positive numbers $\tau\ll t_{0}<T\ll a$ and $a_{1}\ll a$, there exists $C>0$ and $\nu\in(0,1)$ such that, for any smooth section $\sigma$ of $E$ over $X\times\mathbb{R}_{\geq 0}$, we have
$\|\sigma\|_{H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))}\leq C\|\sigma\|_{H_{1}(X\times[0,T])}^{\nu}\Big{(}\|Q\sigma\|_{L^{2}(X\times[0,T])}+\Big{\|}\frac{\partial\sigma}{\partial t}\Big{\|}_{L^{2}(N_{a}(Y)\times\{0\})}\Big{)}^{1-\nu}.$
Proof.
For any $b>0$, we define
$\Omega_{b}=\{(x,t)\in N_{a}(Y)\times[0,a):\varphi_{Y}(x,t)\geq b\}.$
Let $b_{1}<b_{3}<0$ such that
$H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))\subset\Omega_{b_{1}}\subset\Omega_{b_{3}}\subset X\times[0,T].$
We shall prove that there exists a constant $C>0$,
$\|\sigma\|_{H_{1}(\Omega_{b_{1}})}\leq C\|\sigma\|_{H_{1}(\Omega_{b_{3}})}^{\nu}\Big{(}\|Q\sigma\|_{L^{2}(\Omega_{b_{3}})}+\Big{\|}\frac{\partial}{\partial t}\sigma\Big{\|}_{L^{2}(N_{a}(Y)\times\{0\})}\Big{)}^{1-\nu}.$ (5.5)
Let $\rho$ be a smooth non-increasing function on $\mathbb{R}$ such that $\rho(s)=1$ if $s\leq b_{1}$, and $\rho(s)=0$ if $s\geq b_{3}$. Set $\chi=\rho\circ\varphi_{Y}$. It is straightforward that $\nabla\chi$ is only supported on $\Omega_{b_{3}}-\Omega_{b_{1}}$. We fix $b_{2}\in({b_{1}},{b_{3}})$ such that $\rho(s)=1/2$. Let $g=\chi\sigma$, which lies in $\mathcal{F}_{Y}$ by assumption.
First, we consider the right-hand side of line (5.3). Since $\frac{\partial}{\partial t}(\chi\sigma)=\chi\frac{\partial}{\partial t}\sigma+\frac{\partial\chi}{\partial t}\sigma$ and $\frac{\partial\chi}{\partial t}$ is only supported on $\Omega_{b_{3}}-\Omega_{b_{1}}$, we have
$\int_{X\times\mathbb{R}_{\geq 0}}\frac{1}{h}e^{2\varphi_{Y}/h}\Big{|}\frac{\partial}{\partial t}(\chi\sigma)\Big{|}^{2}\geq\frac{1}{8h}e^{2b_{2}/h}\int_{\Omega_{b_{2}}}\Big{|}\frac{\partial}{\partial t}\sigma\Big{|}^{2}-\|\nabla\chi\|_{\infty}\frac{1}{h}e^{2{b_{1}}/h}\int_{\Omega_{b_{3}}-\Omega_{b_{1}}}|\sigma|^{2}.$
Similarly,
$\int_{X\times\mathbb{R}_{\geq 0}}\frac{1}{h}e^{2\varphi_{Y}/h}|\nabla(\chi\sigma)|^{2}\geq\frac{1}{8h}e^{2b_{2}/h}\int_{\Omega_{b_{2}}}|\nabla\sigma|^{2}-\|\nabla\chi\|_{\infty}\frac{1}{h}e^{2{b_{1}}/h}\int_{\Omega_{b_{3}}-\Omega_{b_{1}}}|\sigma|^{2}.$
It is also clear that
$\int_{X\times\mathbb{R}_{\geq 0}}\frac{1}{h^{3}}e^{2\varphi_{Y}/h}|\chi\sigma|^{2}\geq\frac{1}{4h^{3}}e^{2b_{2}/h}\int_{\Omega_{b_{2}}}|\sigma|^{2}.$
Secondly, we consider the left-hand side of line (5.3). It is clear that
$\frac{1}{h}\int_{X\times\{0\}}\Big{|}\frac{\partial}{\partial t}(\chi\sigma)\Big{|}^{2}\leq\frac{1}{h}\int_{N_{b_{3}}(Y)\times\{0\}}\Big{|}\frac{\partial}{\partial t}\sigma\Big{|}^{2}.$
We note that $Q(\chi\sigma)=\chi(Q\sigma)+[Q,\chi]\sigma$, where $[Q,\chi]$ is a first-order differential operator that is supported on $\Omega_{b_{3}}-\Omega_{b_{1}}$. Therefore, there exists $c_{1}>0$ such that
$\begin{split}\int_{X\times\mathbb{R}_{\geq 0}}e^{2\varphi_{Y}/h}|Q(\chi\sigma)|^{2}\leq&\frac{1}{2}\int_{X\times\mathbb{R}_{\geq 0}}e^{2\varphi_{Y}/h}|\chi Q\sigma|^{2}+\int_{\Omega_{b_{3}}-\Omega_{{b_{1}}}}e^{2\varphi_{Y}/h}|[Q,\chi]\sigma|^{2}\\ \leq&\frac{1}{2}\|Q\sigma\|_{L^{2}(\Omega_{b_{3}})}+c_{1}e^{2{b_{1}}/h}\|\sigma\|^{2}_{H^{1}(\Omega_{b_{3}})}.\end{split}$
Combining all the inequalities above, we reach that there exists $c_{2}>0$ such that
$e^{2b_{2}/h}\|\sigma\|_{H^{1}(\Omega_{b_{2}})}^{2}\leq c_{2}e^{2{b_{1}}/h}\|\sigma\|_{H^{1}(\Omega_{b_{3}})}^{2}+c_{2}\Big{(}\|Q\sigma\|_{L^{2}(\Omega_{b_{3}})}+\Big{\|}\frac{\partial}{\partial t}\sigma\Big{\|}_{L^{2}(N_{a}(Y)\times\{0\})}\Big{)}.$
Thus,
$\|\sigma\|_{H^{1}(\Omega_{{b_{1}}})}^{2}\leq c_{2}e^{2({b_{1}}-b_{2})/h}\|\sigma\|_{H^{1}(\Omega_{b_{3}})}^{2}+c_{2}e^{-2b_{2}/h}\Big{(}\|Q\sigma\|_{L^{2}(\Omega_{b_{3}})}+\Big{\|}\frac{\partial}{\partial t}\sigma\Big{\|}_{L^{2}(N_{a}(Y)\times\{0\})}\Big{)}.$
In particular, there exists $h_{0}>0$ such that the above inequality holds uniformly for any $h\in(0,h_{0})$. We emphasize that here ${b_{1}}-b_{2}<0$ and $-b_{2}>0$. Clearly $\|\sigma\|_{H^{1}(\Omega_{{b_{1}}})}\leq\|\sigma\|_{H^{1}(\Omega_{{b_{3}}})}$. This finishes the proof by applying Lemma 5.6. ∎
Lemma 5.8 shows that the $H^{1}$-norm of $\sigma$ on $H^{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))$ is bounded in the sense of interpolation. By the assumption in Definition 4.1, the $r_{2}$-neighborhood of $Y$ covers the entire $X$ for some $r_{2}>0$. We shall prove that the $H^{1}$-norm of $\sigma$ on $X\times N_{\tau}(t_{0})$ is also bounded in the sense of interpolation, by increasing the radius $a_{1}$.
Proposition 5.9.
Let $Y$ be a discrete net in $X$ and $a$ a small positive number. Given small positive numbers $\tau\ll t_{0}<T\ll a$, there exists $C>0$ and $\nu\in(0,1)$ such that for any smooth section $\sigma$ of $E$ over $X\times\mathbb{R}_{\geq 0}$, we have
$\begin{split}&\|\sigma\|_{H_{1}(X\times N_{\tau}(t_{0}))}\\ \leq&C\|\sigma\|_{H_{1}(X\times[0,T])}^{\nu}\Big{(}\|Q\sigma\|_{L^{2}(X\times[0,T])}+\Big{\|}\frac{\partial}{\partial t}\sigma\Big{\|}_{L^{2}(N_{a}(Y)\times\{0\})}\Big{)}^{1-\nu}.\end{split}$ (5.6)
Proof.
We shall prove that there exists $\varepsilon>0$ such that
$\begin{split}&\|\sigma\|_{H_{1}(N_{a_{1}+\varepsilon}(Y)\times N_{\tau}(t_{0}))}\\ \leq&C_{1}\|\sigma\|_{H_{1}(X\times[0,T])}^{\nu_{1}}\Big{(}\|Q\sigma\|_{L^{2}(X\times[0,T])}+\|\sigma\|_{H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))}\Big{)}^{1-\nu_{1}}.\end{split}$ (5.7)
for some $C_{1}>0$ and $\nu_{1}>0$. By Definition 4.1, there are only finitely many steps to exhaust $X$ from $Y$ by increasing $\varepsilon$ of the neighborhood of $Y$. Since $X$ has bounded geometry, Lemma 5.8, Lemma 5.7 and line (5.7) together implies line (5.6).
$Z_{i}$$\varphi_{i}=b_{2}$$\varphi_{i}=b_{1}$$\widehat{\Omega}_{\text{out}}$$z_{i}$$\widehat{\Omega}_{\text{in}}$
We shall prove line (5.7) by applying Lemma 5.5 with carefully chosen $Z$ and $\varphi_{Z}$. Once chosen, the rest of the proof is completely similar to the proof of Lemma 5.6. Given $N_{a_{1}}(Y)=\cup_{i}N_{a_{1}}(y_{i})$, let $z_{i}$ be a point on $\partial N_{a_{1}}(y_{i})$, $Z_{i}$ a tiny piece of $\partial N_{a_{1}}(y_{i})$ near $z_{i}$, and $Z=\cup_{i}Z_{i}$. Pick smooth functions $v_{i}$ supported near $Z_{i}$ such that $|\nabla v_{i}|=1$, $Z_{i}$ is the level set $\{v_{i}=0\}$, and $\nabla v_{i}$ is pointing outward from $Z_{i}\subset\partial N_{a_{1}}(Y)$. We consider the function $\varphi_{Z}$ on $X\times\mathbb{R}_{\geq 0}$ such that
$\varphi_{Z}(x,t)=-v_{i}-d((x,t),(z_{i},t_{0}))^{6}$
near each $Z_{i}$.
Let $\chi$ be a smooth cut-off function that is equal to $1$ on $N_{\varepsilon_{1}}(Z\times\{t_{0}\})$ and equal to 00 outside $N_{\varepsilon_{2}}(Z\times\{t_{0}\})$. Denote $\widehat{\Omega}=N_{\varepsilon_{2}}(Z\times\{t_{0}\})-N_{\varepsilon_{1}}(Z\times\{t_{0}\})$, which contains the supported of $\nabla\chi$. Similar to the proof of Lemma 5.6, we define
$\Omega_{b}=\{(x,t)\in N_{\varepsilon_{1}}(Z\times\{t_{0}\}):\varphi_{Z}(x,t)\geq b\}.$
By construction of $\varphi_{Z}$, we have $\varphi_{Z}(z_{i},t_{0})=0$, and there exists $b_{1}<0$ such that $\Omega_{b_{1}}\cap\widehat{\Omega}$ is contained inside $N_{a_{1}}(Y)\times N_{\tau}(t_{0})$. Pick $b_{2},b_{3}$ with $b_{1}<b_{2}<0<b_{3}$ and $\varepsilon>0$ such that $N_{\varepsilon}(Z)\times N_{\tau}(t_{0})\subset\Omega_{b_{2}}$ and $N_{\varepsilon_{2}}(Z\times\{t_{0}\})\subset(\Omega_{b_{3}})^{c}$.
Now we consider the section $\chi\sigma$, which lies in $\mathcal{F}_{Z}$ by assumption hence satisfies the inequality in Lemma 5.5. Similar to the computation in the proof of Lemma 5.8, the right-hand side of line (5.4) satisfies the following:
$\int_{X\times\mathbb{R}_{\geq 0}}\frac{1}{h}e^{2\varphi_{Z}/h}\Big{(}|\frac{\partial}{\partial t}(\chi\sigma)|^{2}+|\nabla(\chi\sigma)|^{2}\Big{)}\geq$ $\int_{N_{\varepsilon}(Z)\times N_{\tau}(t_{0})}\frac{1}{h}e^{2\varphi_{Z}/h}\Big{(}|\frac{\partial}{\partial t}(\chi\sigma)|^{2}+|\nabla(\chi\sigma)|^{2}\Big{)}$
$\geq$ $\frac{1}{h}e^{2b_{2}/h}\int_{N_{\varepsilon}(Z)\times N_{\tau}(t_{0})}\Big{(}|\frac{\partial}{\partial t}(\chi\sigma)|^{2}+|\nabla(\chi\sigma)|^{2}\Big{)},$
and
$\int_{X\times\mathbb{R}_{\geq 0}}\frac{1}{h^{3}}e^{2\varphi_{Z}/h}|\chi\sigma|^{2}\geq\frac{1}{h^{3}}e^{2b_{2}/h}\int_{N_{\varepsilon}(Z)\times N_{\tau}(t_{0})}|\sigma|^{2}.$
For the left-hand side of line (5.4), we still notice that
$Q(\chi\sigma)=\chi Q\sigma+[Q,\chi]\sigma,$
where $[Q,\chi]$ is a first-order differential operator that is supported only on $\widehat{\Omega}$. Write $\widehat{\Omega}=\widehat{\Omega}_{\text{in}}\cup\widehat{\Omega}_{\text{out}}$, where $\widehat{\Omega}_{\text{in}}\coloneqq\Omega_{b_{1}}\cap\widehat{\Omega}$ and $\widehat{\Omega}_{\text{out}}$ is the complement of $\widehat{\Omega}_{\text{in}}$. We note that by construction, $\widehat{\Omega}_{\text{in}}$ is contained in $N_{a_{1}}(Y)\times N_{\tau}(t_{0})$, while on $\widehat{\Omega}_{\text{out}}$ we have $\varphi_{Z}\leq b_{1}$. Therefore, there exists $c_{1}>0$ such that
$\int_{X\times\mathbb{R}_{\geq 0}}e^{2\varphi_{Z}/h}|Qg|^{2}\leq\frac{1}{2}\int_{N_{\varepsilon_{1}}(Z\times\{t_{0}\})}e^{2\varphi_{Z}/h}|\chi Q\sigma|^{2}+\int_{\widehat{\Omega}}e^{2\varphi_{Z}/h}|[Q,\chi]\sigma|^{2}$
$\leq$ $\frac{1}{2}e^{2b_{3}/h}\int_{L^{2}(N_{\varepsilon_{1}}(Z\times\{t_{0}\}))}|Q\sigma|^{2}+\int_{\widehat{\Omega}_{\text{in}}}e^{2\varphi_{Z}/h}|[Q,\chi]\sigma|^{2}+\int_{\widehat{\Omega}_{\text{out}}}e^{2\varphi_{Z}/h}|[Q,\chi]\sigma|^{2}$
$\leq$ $\frac{1}{2}e^{2b_{3}/h}\|Q\sigma\|^{2}_{L^{2}(X\times[0,T])}+c_{1}e^{2b_{3}/h}\|\sigma\|^{2}_{H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))}+c_{1}e^{2b_{1}/h}\|\sigma\|^{2}_{H^{1}(X\times[0,T])}$
Therefore, by Lemma 5.5, there is $c_{2}>0$ such that
$e^{2b_{2}/h}\|\sigma\|_{H^{1}(N_{\varepsilon}(Z)\times N_{\tau}(t_{0}))}^{2}$
$\leq$ $c_{2}\Big{(}e^{2b_{3}/h}\|Q\sigma\|^{2}_{L^{2}(X\times[0,T])}+e^{2b_{3}/h}\|\sigma\|^{2}_{H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))}+e^{2b_{1}/h}\|\sigma\|^{2}_{H^{1}(X\times[0,T])}\Big{)}.$
Equivalently,
$\|\sigma\|_{H^{1}(N_{\varepsilon}(Z)\times N_{\tau}(t_{0}))}^{2}$
$\leq$ $c_{2}e^{2(b_{1}-b_{2})/h}\|\sigma\|^{2}_{H^{1}(X\times[0,T])}+c_{2}e^{2(b_{3}-b_{2})/h}\Big{(}\|Q\sigma\|^{2}_{L^{2}(X\times[0,T])}+\|\sigma\|^{2}_{H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))}\Big{)}.$
We note that $b_{1}-b_{2}<0$ and $b_{3}-b_{2}>0$. It follows together with Lemma 5.6 that
$\|\sigma\|_{H^{1}(N_{\varepsilon}(Z)\times N_{\tau}(t_{0}))}\leq c_{3}\|\sigma\|^{\nu_{1}}_{H^{1}(X\times[0,T])}\Big{(}\|Q\sigma\|^{2}_{L^{2}(X\times[0,T])}+\|\sigma\|^{2}_{H_{1}(N_{a_{1}}(Y)\times N_{\tau}(t_{0}))}\Big{)}^{1-\nu_{1}}.$
for some $c_{3}>0$ and $\nu_{1}>0$. Note that as $X$ has bounded geometry $N_{a_{1}+\varepsilon}(Y)$ is covered by at most $N$ sets, which are of the form $N_{\varepsilon}(Z)$ for some $Z\subset\partial N_{a_{1}}(Y)$. This finishes the proof of line (5.7) with $C_{3}=Nc_{1}$, hence complete the proof of line (5.6) by the discussion at the beginning. ∎
Finally, we are ready to prove the Theorem 5.1.
Proof.
The Gårding inequality implies that
$\langle P\psi,\psi\rangle\geq c\|\nabla\psi\|^{2}-c^{\prime}\|\psi\|^{2},$ (5.8)
for any $L^{2}$-section $\psi$. Thus, we obtain that $P+c^{\prime}\geq 0$. Without loss of generality, we may assume that $P\geq 0$.
Given $\sigma\in V_{\lambda}$, we define
$F_{t}=\frac{\sinh(t\sqrt{P})}{\sqrt{P}}\sigma=\frac{e^{t\sqrt{P}}-e^{-t\sqrt{P}}}{2\sqrt{P}}\sigma=\sum_{n=0}^{\infty}\frac{t^{2n+1}}{(2n+1)!}P^{n}\sigma.$
It is clear from the definition that
$\frac{\partial}{\partial t}F_{t}\Big{|}_{t=0}=\sigma,\text{ and }QF_{t}=\Big{(}-\frac{\partial^{2}}{\partial t^{2}}+P\Big{)}F_{t}=0.$
Together with the Proposition 5.9, we obtain that
$\|F_{t}\|_{H_{1}(X\times(t_{0}-\tau,t_{0}+\tau))}\leq C\|F_{t}\|_{H_{1}(X\times[0,T])}^{\nu}\|\sigma\|_{L^{2}(N_{a}(Y))}^{1-\nu}.$ (5.9)
Moreover, the construction of $F_{t}$ directly implies that, for any $t\in[0,T]$,
$\|F_{t}\|\leq\frac{\sinh(t\sqrt{\lambda})}{\sqrt{\lambda}}\|\sigma\|,~{}\langle PF_{t},F_{t}\rangle\leq\sqrt{\lambda}\sinh(t\sqrt{\lambda})\|\sigma\|,$
and
$\|\sigma\|\leq\Big{\|}\frac{\partial}{\partial t}F_{t}\Big{\|}\leq\cosh(t\sqrt{\lambda})\|\sigma\|.$
Therefore,
[…]
$\leq$ $T\Big{(}(1+\frac{c^{\prime}}{c}\frac{\sinh^{2}(T\sqrt{\lambda})}{\lambda}+\cosh^{2}(T\sqrt{\lambda}))+\frac{1}{c}\lambda\sinh^{2}(T\lambda)\Big{)}\|\sigma\|^{2}$
$=$ $C_{\lambda}^{\frac{2(1-\nu)}{\nu}}\big{(}\frac{2\tau}{C^{2}}\big{)}^{\frac{1}{\nu}}\|\sigma\|^{2}$
Here,
$C_{\lambda}=\big{(}\frac{C^{2}}{2\tau}\big{)}^{\frac{1}{2(1-\nu)}}T^{\frac{\nu}{2(1-\nu)}}\Big{(}(1+\frac{c^{\prime}}{c}\frac{\sinh^{2}(T\sqrt{\lambda})}{\lambda}+\cosh^{2}(T\sqrt{\lambda}))+\frac{1}{c}\lambda\sinh^{2}(T\lambda)\Big{)}^{\frac{\nu}{2(1-\nu)}}.$
By the Gårding inequality, we have that
[…]
Thus we obtain from line (5.9) that
$2\tau\|\sigma\|^{2}\leq C^{2}\Big{(}C_{\lambda}^{\frac{2(1-\nu)}{\nu}}\big{(}\frac{2\tau}{C^{2}}\big{)}^{\frac{1}{\nu}}\|\sigma\|^{2}\Big{)}^{\nu}\|\sigma\|_{L^{2}(N_{a}(Y))}^{2-2\nu}.$
A direct simplification indicates that
$\|\sigma\|_{L^{2}(X)}\leq C_{\lambda}\|\sigma\|_{L^{2}(N_{a}(Y))}.$
Appendix A Invertible doubles
In the appendix, we prove an invertibility theorem on doubles, as an application of Theorem 5.1, which is of independent interest.
In [Xieyu14], Xie and Yu proved a higher analogue of the relative index theorem. A key ingredient is the vanishing theorem of the index of an elliptic differential operator on the double of a complete manifold, which is proved in [Xieyu14, Remark 5.5]. In this section, we prove a stronger result that the elliptic operator on a double is indeed invertible, based on Theorem 5.1.
For simplicity, we only consider the invertibility of Dirac operator on the double. Let $(M_{1},g)$ be an even dimensional closed spin Riemannian manifold with boundary $N=\partial M_{1}$, where $N$ is a closed manifold. We denote a copy of $M_{1}$ with the reserved orientation by $M_{2}=-M_{1}$. Let $(\widetilde{M}_{1},\widetilde{g})$ be a regular $\Gamma$-cover of $(M_{1},g)$ equipped with the lifted metric, and $(\widetilde{M}_{2},\widetilde{g})$ the $\Gamma$-cover of $(M_{2},g)$. We glue $\widetilde{M}_{1}$ and $\widetilde{M}_{2}$ along a tubular neighborhood of the boundary and obtain a double $\widetilde{M}$ of $\widetilde{M}_{1}$. Denote by $S_{\widetilde{M}_{1}}$, $S_{\widetilde{M}_{2}}$ and $S_{\widetilde{M}}$ the spinor bundles over $\widetilde{M}_{1}$, $\widetilde{M}_{2}$ and $\widetilde{M}_{1}$, respectively. Now the spinor bundles are glued together by the Clifford action $c(v)$, where $v$ is the inward unit normal vector near the boundary of $Y_{1}$. Note that
$S_{\widetilde{M}}^{\pm}=S_{\widetilde{M}_{1}}^{\pm}\cup_{c(v)}S_{\widetilde{M}_{2}}^{\mp}$
In particular, a section of $S_{\widetilde{M}}$ can be identified with a pair $(\sigma_{1},\sigma_{2})$ such that $\sigma_{1}$ is a section of $S_{\widetilde{M}_{1}}$, $\sigma_{2}$ is a section of $S_{\widetilde{M}_{2}}$, and on the boundary $\widetilde{N}\coloneqq\partial\widetilde{M}_{1}$
$\sigma_{2}=c(v)\sigma_{1}.$
The Dirac operator $\widetilde{D}_{i}$ acting on $S_{\widetilde{M}_{i}}$ is an odd operator given by $\widetilde{D}_{i}^{\pm}\colon S_{\widetilde{M}_{i}}^{\pm}\to S_{\widetilde{M}_{i}}^{\mp}$ , and the Dirac operator $\widetilde{D}$ acting on $S_{\widetilde{M}}$ is identified with
$\widetilde{D}^{\pm}(\sigma_{1},\sigma_{2})=(\widetilde{D}_{1}^{\pm}\sigma_{1},\widetilde{D}_{2}^{\mp}\sigma_{2}).$
The main result of the appendix is as follows.
Theorem A.1.
The Dirac operator $\widetilde{D}$ on $\widetilde{M}$ is invertible, i.e., there exists $C>0$ such that
$\|\sigma\|\leq C\|\widetilde{D}\sigma\|$
for any smooth $L^{2}$-section $\sigma$ of $S_{\widetilde{M}}$. In particular, the higher index $\textup{Ind}(\widetilde{D})\in K_{*}(C^{*}_{r}(\Gamma))$ is zero.
We first construct extension maps by the following lemma.
Lemma A.2.
With the notation above, there exist bounded linear maps
$\mathscr{E}_{i}\colon L^{2}(\widetilde{N},S_{\widetilde{M}})\to H^{1}(\widetilde{M}_{i},S_{\widetilde{M}_{i}}),$
for $i=1,2$, such that for any $\psi\in L^{2}(\widetilde{N})$
$\psi=(\mathscr{E}_{1}\psi)|_{\widetilde{N}}-c(v)\cdot(\mathscr{E}_{2}\psi)|_{\widetilde{N}}.$
Proof.
Let $\Omega$ be a tubular neighborhood $[0,\tau]\times\widetilde{N}$ near $\widetilde{N}$ for $\tau$ small enough. Equip $\Omega$ with the product metric $dt^{2}+g_{\widetilde{N}}$. Since the lifted metrics have bounded geometry, the $H^{1}$-norms of $S_{\widetilde{M}_{i}}$ over $\Omega$ induced by the product metric and the metric $\widetilde{g}$ are equivalent. We will show the boundedness of $\mathscr{E}_{i}$ using the product metric.
Let $\widehat{D}$ be the Dirac operator on $\Omega$ with respect to the product metric, namely
$\widehat{D}=c(v)\frac{\partial}{\partial t}+\sum_{i=1}^{n-1}c(e_{i})\nabla_{e_{i}}^{\widetilde{N}},$
where $e_{i}$’s are local orthonormal basis of $\widetilde{N}$ and $\nabla^{\widetilde{N}}$ the spinorial connection on $\widetilde{N}$. Set
$\widehat{D}_{\partial}\coloneqq-c(v)\sum_{i=1}^{n-1}c(e_{i})\nabla_{e_{i}}^{\widetilde{N}},$
which is a self-adjoint operator acting on $S_{\widetilde{M}}$ over $\widetilde{N}$. Let $P_{+}$ be the spectral projection of the non-negative spectrum of $\widehat{D}_{\partial}$, and $P_{-}$ be the spectral projection of the negative spectrum of $\widehat{D}_{\partial}$.
Given $\psi\in L^{2}(\widetilde{N})$, consider the section $e^{-t\widehat{D}_{\partial}}P_{+}\psi$ for $t\in[0,\tau]$. We see that
[…]
and
$\widehat{D}(e^{-t\widehat{D}_{\partial}}P_{+}\psi)=c(v)\Big{(}\frac{\partial}{\partial t}+\widehat{D}_{\partial}\Big{)}e^{-t\widehat{D}_{\partial}}P_{+}\psi=0.$
Therefore, if we pick a cut-off function $\chi(t)$ that is equal to $1$ near $t=0$ and supported in $[0,\tau)$, we see that $\chi(t)e^{-t\widehat{D}_{\partial}}P_{+}\psi$ is an $H^{1}$-section, and the map
$\mathscr{E}_{1}\colon\psi\mapsto\chi(t)e^{-t\widehat{D}_{\partial}}P_{+}\psi$
is a bounded map from $L^{2}(\widetilde{N},S_{\widetilde{M}})$ to $H^{1}(\widetilde{M}_{1},S_{\widetilde{M}_{1}})$.
Note that $c(v)$ anti-commutes with $\widehat{D}_{\partial}$, hence $c(v)P_{-}\psi$ lie the range of the positive spectrum of $\widehat{D}_{\partial}$. Therefore, the map
$\mathscr{E}_{2}\colon\psi\mapsto\chi(t)e^{-t\widehat{D}_{\partial}}c(v)P_{-}\psi$
is a bounded map from $L^{2}(\widetilde{N},S_{\widetilde{M}})$ to $H^{1}(\widetilde{M}_{2},S_{\widetilde{M}_{2}})$. Clearly, we have
$\psi=P_{+}\psi+P_{-}\psi=P_{+}\psi-c(v)\cdot\big{(}c(v)P_{-}\psi\big{)}=(\mathscr{E}_{1}\psi)|_{\widetilde{N}}-c(v)\cdot(\mathscr{E}_{2}\psi)|_{\widetilde{N}}.$
This finishes the proof. ∎
Proof of Theorem A.1.
Assume on the contrary that $\widetilde{D}$ is not invertible. Therefore, without loss of generality, for any $\varepsilon>0$, there exists a smooth $L^{2}$-section $\sigma=(\sigma_{1},\sigma_{2})$ of $S_{\widetilde{M}}^{+}$, such that $\|\sigma\|=1$ and $\|\widetilde{D}\sigma\|<\varepsilon$. We assume that $\varepsilon\leq 1$. As $\|\sigma\|^{2}=\|\sigma_{1}\|^{2}+\|\sigma_{2}\|^{2}$, we may assume that $\|\sigma_{1}\|\geq\sqrt{2}/2$. Note that
$\|\widetilde{D}\sigma\|=\|\widetilde{D}^{+}\sigma\|=\|(\widetilde{D}_{1}^{+}\sigma_{1},\widetilde{D}_{2}^{-}\sigma_{2})\|=\sqrt{\|\widetilde{D}_{1}^{+}\sigma_{1}\|^{2}+\|\widetilde{D}_{2}^{-}\sigma_{2}\|^{2}}\leq\varepsilon.$
Thus $\|\widetilde{D}_{1}\sigma_{1}\|<\varepsilon$.
On the boundary $\widetilde{N}$, we have $\sigma_{2}=c(v)\sigma_{1}$ by assumption. Therefore, by the divergence theorem, we have
$\int_{\widetilde{M}_{1}}\langle\widetilde{D}^{+}\sigma_{1},\sigma_{2}\rangle-\int_{\widetilde{M}_{1}}\langle\sigma_{1},\widetilde{D}^{-}\sigma_{2}\rangle=-\int_{\widetilde{N}}\langle c(v)\sigma_{1},\sigma_{2}\rangle=\int_{\widetilde{N}}|\sigma_{1}|^{2}.$
It follows that
$\|\sigma_{1}\|_{L^{2}(\widetilde{N})}^{2}\leq 2\|\widetilde{D}\sigma\|\|\sigma\|\leq 2\varepsilon.$
For $i=1,2$, let
$\mathscr{E}_{i}\colon L^{2}(\widetilde{N},S_{\widetilde{M}})\to H^{1}(\widetilde{M}_{i},S_{\widetilde{M}_{i}}),$
be the extension maps constructed in Lemma A.2. Suppose that $\|\mathscr{E}_{i}\|\leq C_{1}$ for some $C_{1}>0$. Write for short $\psi=\sigma_{1}|_{\widetilde{N}}$. We define
$\sigma^{\prime}=(\sigma^{\prime}_{1},\sigma^{\prime}_{2})\coloneqq(\sigma_{1}-\mathscr{E}_{1}(\psi),\mathscr{E}_{2}(\psi)).$
By Lemma A.2, we have
$\sigma^{\prime}_{2}=E_{2}(\psi)|_{\widetilde{N}}=c(v)\psi-c(v)\cdot\mathscr{E}_{1}(\psi)|_{\widetilde{N}}=c(v)\sigma_{1}^{\prime}.$
Therefore, $\sigma^{\prime}$ is a well-define section of $H^{1}(\widetilde{M},S_{\widetilde{M}})$. Furthermore, we have
$\|\widetilde{D}\sigma^{\prime}\|\leq\|\widetilde{D}_{1}\sigma_{1}\|+n\|\mathscr{E}_{1}(\psi)\|_{H^{1}}+n\|\mathscr{E}_{2}(\psi)\|_{H^{1}}\leq\varepsilon+2C_{1}\sqrt{2\varepsilon}\leq C_{2}\sqrt{\varepsilon}$
for some $C_{2}>0$. Since $\|\sigma_{1}\|\geq\sqrt{2}/2$, we have
$\|\sigma_{1}^{\prime}\|\geq\sqrt{2}/2-C_{1}\sqrt{2\varepsilon},\text{ and }\|\sigma_{2}^{\prime}\|\leq C_{1}\sqrt{2\varepsilon}.$
Let $P_{\sqrt{\varepsilon}}$ be the spectral projection of $\widetilde{D}^{2}$ with spectrum $\leq\sqrt{\varepsilon}$. Write
$\sigma^{\prime}=P_{\sqrt{\varepsilon}}\sigma^{\prime}+(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}.$
Clearly $P_{\sqrt{\varepsilon}}\sigma^{\prime}\perp(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}$ and $\widetilde{D}P_{\sqrt{\varepsilon}}\sigma^{\prime}\perp\widetilde{D}(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}$. As $(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}$ lies in the range where the spectrum of $\widetilde{D}^{2}$ is $\geq\sqrt{\varepsilon}$, we have
$\|\widetilde{D}(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}\|\geq\varepsilon^{1/4}\|(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}\|.$
Therefore
$\|(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}\|\leq\varepsilon^{-1/4}\|\widetilde{D}(1-P_{\sqrt{\varepsilon}})\sigma^{\prime}\|\leq\varepsilon^{-1/4}\|\widetilde{D}\sigma^{\prime}\|\leq C_{2}\varepsilon^{1/4}.$
Set
$\sigma^{\prime\prime}=P_{\sqrt{\varepsilon}}\sigma^{\prime}=(\sigma^{\prime\prime}_{1},\sigma^{\prime\prime}_{2}).$
Since we have shown that $\|\sigma^{\prime}-\sigma^{\prime\prime}\|\leq C_{2}\varepsilon^{1/4}$, we see that
$\|\sigma_{1}^{\prime\prime}\|\geq\sqrt{2}/2-C_{1}\sqrt{2\varepsilon}-C_{2}\varepsilon^{1/4},\text{ and }\|\sigma_{2}^{\prime}\|\leq C_{1}\sqrt{2\varepsilon}+C_{2}\varepsilon^{1/4}.$
In particular, there is $C_{3}>0$ such that
$\|\sigma_{2}^{\prime\prime}\|\leq C_{3}\varepsilon^{1/4}\|\sigma^{\prime\prime}\|.$
However, since $\widetilde{M}_{2}$ contains an $a$-neighborhood of a discrete net in $\widetilde{M}$, by Theorem 5.1, there exists $C_{4}>0$ such that
$\|\sigma^{\prime\prime}\|\leq C_{4}\|\sigma_{2}^{\prime\prime}\|$
uniformly for any $\varepsilon\leq 1$. This leads to a contradiction as $\|\sigma^{\prime\prime}\|\geq\|\sigma_{1}^{\prime\prime}\|$, which is bounded away from zero. ∎